### v1

# The quantum-refractive evolution of polarization states in pulsar emission

Abstract
Highly magnetized neutron stars have quantum refraction effects on pulsar emission due to the non-linearity of the quantum electrodynamics (QED) action. In this paper, we investigate the evolution of the polarization states under the quantum refraction effects combined with the frequency dependence of pulsar emission; we solve a system of evolution equations of the Stokes vector, where the birefringent vector, in which such effects are encoded, acts on the Stokes vector. At a fixed frequency of emission, depending on the magnitude of the birefringent vector, dominated mostly by the magnetic field strength, the evolution of the Stokes vector largely exhibits three different patterns: (i) monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviours. These features are understood and confirmed by means of approximate analytical solutions to the evolution equations.
keywords:
pulsars – magnetic fields – photon emission – non-linear electrodynamics – quantum refraction – polarization evolution
pubyear: 2024pagerange: The quantum-refractive evolution of polarization states in pulsar emissionB
1 Introduction
Highly magnetized neutron stars are interesting astrophysical objects that will provide an astrophysical arena for probing fundamental physics (Raffelt (1996)), in particular, strong field QED (Meszaros (1992); Harding & Lai (2006)). The magnetic fields of neutron stars of $10^{12}-10^{15}\,\mathrm{G}$ are strong enough to give rise to QED effects, such as vacuum polarization and pair production due to the induced electric fields, which have been proposed as an astrophysical laboratory to test strong field QED (see Kim & Kim (2021) and references therein). A great advantage of neutron stars as a laboratory is virtually infinitely long duration and range of the electromagnetic fields (in the scale of Compton time and length), strongly contrasting with the short, pulsed fields achieved with ultra-intense lasers, as well as the strength of the fields (Yoon et al. (2021)).
Strong electromagnetic fields polarize the vacuum, which leads to the Heisenberg-Euler-Schwinger (HES) one-loop effective action (Heisenberg & Euler (1936); Schwinger (1951)). The effective action describes non-linear electrodynamics, which has an imaginary part from the vacuum instability due to production of charged particle pairs via electric fields. The HES action leads to the vacuum multiple refraction indices, which is generally referred to as the quantum refraction effect. A test of strong-field QED is one of the most important problems in physics, and there have been laboratory experiment proposals for it, e.g., PVLAS (Polarizzazione del Vuoto con LAser) project (Ejlli et al. (2020)). On the other hand, highly magnetized neutron stars provide near-critical fields, and these can be used as various celestial laboratories to test strong field QED (Ruffini et al. (2010); Fedotov et al. (2023); Hattori et al. (2023))). Also, there have been space missions proposed to observe the strong-field QED effects, e.g., the enhanced X-ray Timing and Polarimetry (eXTP) (Santangelo et al. (2019)) and the Compton Telescope project (Wadiasingh et al. (2019)).
The HES action is well approximated by the post-Maxwellian action, even up to the strength one order lower than the critical magnetic field $B_{\mathrm{c}}=m_{\mathrm{e}}^{2}c^{3}/e\hbar=4.4\times 10^{13}\,\mathrm{G}$, which keeps up to the quadratic terms of the Maxwell scalar and pseudo-scalar. Therefore, the post-Maxwellian action exhibits non-linear characteristics of vacuum polarization, such as quantum refraction. Previously, we have studied the quantum refraction effects on the polarization vectors in the dipole magnetic field background of a pulsar model (Kim et al. (2024)). The study is non-trivial in comparison with other similar studies where the background magnetic field is assumed to be uniform, in that we have to deal with dipole magnetic field, whose strength and direction vary over space.
In this paper, we study the evolution of the polarization states under the quantum refraction effects combined with the frequency dependence of pulsar emission. Our analysis is based on the evolution equations of the Stokes vector, where such effects are encoded into the birefringent vector that acts on the Stokes vector. The Stokes vector has a crucial advantage over the polarization vector in representing polarization states: it can be directly determined from experimentally measurable quantities and accommodate depolarization effects due to incomplete coherence and random processes during the photon propagation. Solutions of the evolution equations describe how the polarization states change along the photon propagation path from the emission point towards an observer. It turns out that the evolution of the Stokes vector, at a fixed frequency of emission, largely exhibits three different patterns, depending on the magnitudes of the birefringent vector, dominated mostly by the magnetic field strength: (i) fractionally oscillatory - monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviours, which are found by numerical solutions and also confirmed by approximate analytical solutions. These are novel features rarely illuminated in previous studies on the same topic.
The paper is organized as follows. In Section 2.1, we introduce a system of evolution equations of the Stokes vector and apply this formalism to our pulsar emission model for an oblique rotator. In Section 2.2, the evolution equations are solved for some known rotation-powered pulsars (RPPs) in three ways: fully numerically, via perturbation analysis, and using an analytical approximation. Also, we discuss the evolution patterns of the Stokes vectors resulting from the solutions. Then finally, we conclude the paper with discussions on other similar studies and follow-up studies.
2 Evolution of Stokes vector in strong magnetic field
2.1 Evolution equations of Stokes vector
Classically, polarization properties of pulsar emission are described by the Stokes parameters $\left\{I,Q,U,V\right\}$, where $I$ is a measure of the total intensity, $Q$ and $U$ jointly describe the linear polarization, and $V$ describes the circular polarization of the curvature emission in pulsars (for more details, see Appendix A). However, in the presence of strong magnetic fields in the background of the emission, the polarization evolves along the photon propagation path from the emission point towards an observer. The evolution of the polarization can be investigated systematically using the formalism initiated by Kubo & Nagata (1981, 1983, 1985), and further developed by Heyl & Shaviv (2000), Heyl et al. (2003), Heyl & Caiazzo (2018), and Novak et al. (2018), namely, a system of evolution equations of the Stokes vector, described as
$\frac{\mathrm{d}\mathbf{S}}{\mathrm{d}s}=k\mathbf{\hat{\Omega}}\times\mathbf{S},$ (1)
where $k\equiv\omega/c$ denotes the wave number for the electromagnetic radiation and $s$ is an affine parameter to measure the length of the photon trajectory, and $\mathbf{S}$ is the normalized Stokes vector, defined out of the Stokes parameters as $\mathbf{S}=\left(S_{1},S_{2},S_{3}\right)\allowbreak\equiv\left(Q/I,U/I,V/I\right)$, [^1] The classical Stokes vector can be expressed via pulse profiles of pulsar curvature emission, as illustrated in Appendix A. and $\mathbf{\hat{\Omega}}$ is the birefringent vector, defined as [^2] Note that our $k\mathbf{\hat{\Omega}}$ is equivalent to the birefringent vector as defined in the references above.
$\mathbf{\hat{\Omega}}\equiv\frac{\alpha_{\mathrm{e}}}{30\pi}\left(B/B_{\mathrm{c}}\right)^{2}\sin^{2}\vartheta\left(\mathcal{E}_{\mathrm{I}}^{2}-\mathcal{E}_{\mathrm{II}}^{2},2\mathcal{E}_{\mathrm{I}}\mathcal{E}_{\mathrm{II}},0\right),$ (2)
where $\alpha_{\mathrm{e}}$ denotes the fine-structure constant and $\alpha_{\mathrm{e}}/\left(30\pi\right)\approx 7.743\times 10^{-5}$ and $B_{\mathrm{c}}\approx 4.414\times 10^{13}\,\mathrm{G}$ is the critical magnetic field, and $\vartheta$ denotes the angle between the photon trajectory and the local magnetic field line (see Fig. 1), i.e.,
$\vartheta=\cos^{-1}\left(\mathbf{\hat{n}}_{\left[0\right]}\cdot\mathbf{\hat{B}}\right),$ (3)
with $\mathbf{\hat{n}}_{\left[0\right]}$ being the classical propagation vector and $\mathbf{\hat{B}}\equiv\mathbf{B}/\left|\mathbf{B}\right|$, and
$\mathcal{E}_{i}\equiv-\mathbf{\hat{B}\cdot}\left(\mathbf{\hat{n}}_{\left[0\right]}\times\boldsymbol{\varepsilon}_{i\left[0\right]}\right),~{}~{}i=\mathrm{I},\mathrm{II},$ (4)
with $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ being the two classical mode polarization vectors, orthogonal to each other and to $\mathbf{\hat{n}}_{\left[0\right]}$; the specific forms of $\mathbf{\hat{n}}_{\left[0\right]}$, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ are later given by equations (8), (12) and (13), respectively, for the magnetic field of an oblique dipole rotator as described by equation (5) and Fig. 1.
Figure 1: A cross-sectional view of a pulsar magnetosphere with the dipole magnetic field lines (green) around a neutron star. The vertical dashed line (black) and the inclined solid line (red) represent the rotation axis and the magnetic axis, respectively. $\alpha$ between these axes denotes the inclination angle. The scale of the unity in this graph is equivalent to the neutron star radius $\sim 10^{6}\,\mathrm{cm}$. The red dotted line represents the trajectory curve of the light ray traced by $\mathbf{\hat{n}}_{\left[0\right]}$ as projected onto the $xz$-plane. (Credit: Kim & Trippe (2021), reproduced with modifications.)
In our pulsar emission model, we consider curvature radiation produced along the magnetic field lines of an oblique dipole rotator as illustrated in Fig. 1:
$\mathbf{B}\left(r,\theta,\phi\right)=\frac{2\mu\left(\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi\right)}{r^{3}}\mathbf{e}_{\hat{r}}+\frac{\mu\left(\cos\alpha\sin\theta-\sin\alpha\cos\theta\cos\phi\right)}{r^{3}}\mathbf{e}_{\hat{\theta}}+\frac{\mu\sin\alpha\sin\phi\,}{r^{3}}\mathbf{e}_{\hat{\phi}},$ (5)
where $\mu$ is the magnetic dipole moment and $\alpha$ denotes the inclination angle between the rotation axis and the magnetic axis. [^3] Here the symbol $\alpha$ must be distinguished from the fine-structure constant $\alpha_{\mathrm{e}}$. The photon beam from curvature radiation is tangent to the field line at the emission point $\left(x_{\mathrm{o}},y_{\mathrm{o}},z_{\mathrm{o}}\right)=\left(r_{\mathrm{o}}\sin\theta_{\mathrm{o}},0,r_{\mathrm{o}}\cos\theta_{\mathrm{o}}\right)$.
However, at the same time, our pulsar magnetosphere rotates, and therefore the field lines get twisted due to the magneto-centrifugal acceleration on the plasma particles moving along the field lines (Blandford & Payne (1982)). Then, taking into consideration this magneto-hydrodynamic (MHD) effect, the direction of the classical photon propagation, which must line up with the particle velocity in order for an observer to receive the radiation, can be described as (Gangadhara (2005))
$\mathbf{\hat{n}}_{\left[0\right]}=\beta\mathbf{\hat{B}}+\mathbf{\Omega}_{\mathrm{p}}\times\mathbf{r,}$ (6)
where on the right-hand side
$\beta\equiv\left[1-\left(\frac{\Omega_{\mathrm{p}}r}{c}\right)^{2}\sin^{2}\theta\left(1-\frac{\sin^{2}\alpha\sin^{2}\phi}{3\cos^{2}\theta^{\prime}+1}\right)\right]^{1/2}-\frac{\Omega_{\mathrm{p}}r}{c}\frac{\sin\alpha\sin\theta\sin\phi}{\left(3\cos^{2}\theta^{\prime}+1\right)^{1/2}},$ (7)
with $c$ being the speed of light and $\cos\theta^{\prime}\equiv\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi$, and the second term accounts for the centrifugal acceleration, with $\mathbf{\Omega}_{\mathrm{p}}\equiv\Omega_{\mathrm{p}}\mathbf{e}_{z}$ [^4] Here the symbol $\mathbf{\Omega}_{\mathrm{p}}$ must be distinguished from the birefringent vector $\mathbf{\hat{\Omega}}$. and $\Omega_{\mathrm{p}}=2\pi/P$ being a pulsar rotation (angular) frequency, as given in terms of the rotation period $P$.
During the rotation the azimuthal phase changes by $\phi\sim\Omega_{\mathrm{p}}t$, while our photon has propagated a distance by $s\sim ct$. In our analysis, the photon propagation is described with the consideration of the MHD effect above, assuming $\phi$ to be very small; e.g., $\phi\lesssim 10^{-1}$ is considered for a millisecond pulsar with $\Omega_{\mathrm{p}}\sim 10^{2}\,\mathrm{Hz}$, during the time of rotation $t\lesssim 10^{-3}\,\mathrm{s}$, such that $s\lesssim 10^{7}\,\mathrm{cm}$, which corresponds to the propagation distance within about $10$ times the neutron star radius. For equation (6) we take only the leading order expansions of $\mathbf{\hat{B}}\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ and $\beta\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ in $\phi$ from equations (5) and (7), respectively, and can express the classical propagation vector $\mathbf{\hat{n}}_{\left[0\right]}$ in Cartesian coordinates as
$\mathbf{\hat{n}}_{\left[0\right]}=\hat{n}_{x\left[0\right]}\mathbf{e}_{x}+\hat{n}_{y\left[0\right]}\mathbf{e}_{y}+\hat{n}_{z\left[0\right]}\mathbf{e}_{z}\$ (8)
with
$\hat{n}_{x\left[0\right]}$ $\approx\frac{2\cos\left(\theta_{\mathrm{o}}-\alpha\right)\sin\theta_{\mathrm{o}}+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\cos\theta_{\mathrm{o}}}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (9)
$\hat{n}_{z\left[0\right]}$ $\approx\frac{2\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\theta_{\mathrm{o}}-\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\theta_{\mathrm{o}}}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (10)
and
$\hat{n}_{y\left[0\right]}\approx\frac{\Omega_{\mathrm{p}}}{c}\left[\frac{\sin\alpha s}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}+r_{\mathrm{o}}\sin\theta_{\mathrm{o}}\right]+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (11)
where we have considered $\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\lesssim\phi$, e.g., for a millisecond pulsar with $\Omega_{\mathrm{p}}\sim 10^{2}\,\mathrm{Hz}$ and $r_{\mathrm{o}}\sim 10^{6}\,\mathrm{cm}$, such that $\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\lesssim\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\lesssim\phi^{2}$, all to be ignored in our analysis, and have substituted $\phi=\Omega_{\mathrm{p}}s/c$ in equation (11), the leading order rotational effect to be considered in our analysis.
The orthogonal pair of classical mode polarization vectors, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$, both being also orthogonal to $\mathbf{\hat{n}}_{\left[0\right]}$ as given by equation (8) above, are determined as
$\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ $=\hat{n}_{z\left[0\right]}\mathbf{e}_{x}+\hat{n}_{y\left[0\right]}\mathbf{e}_{y}-\hat{n}_{x\left[0\right]}\mathbf{e}_{z},$ (12)
$\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ $=-\left(\hat{n}_{x\left[0\right]}+\hat{n}_{z\left[0\right]}\right)\hat{n}_{y\left[0\right]}\mathbf{e}_{x}+\mathbf{e}_{y}+\left(\hat{n}_{x\left[0\right]}-\hat{n}_{z\left[0\right]}\right)\hat{n}_{y\left[0\right]}\mathbf{e}_{z},$ (13)
such that the three vectors, $\mathbf{\hat{n}}_{\left[0\right]}$, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ form an orthonormal basis. [^5] It can be checked out that $\mathbf{\hat{n}}_{\left[0\right]}\cdot\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}\approx 0+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$, $\mathbf{\hat{n}}_{\left[0\right]}\cdot\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}=0$ and $\boldsymbol{\varepsilon}_{\,\mathrm{I}\left[0\right]}\cdot\boldsymbol{\ \varepsilon}_{\mathrm{II}\left[0\right]}=0$ while $\mathbf{\hat{n}}_{\left[0\right]}^{2}\approx 1+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$ and $\boldsymbol{\ \varepsilon}_{\mathrm{I},\mathrm{II}\left[0\right]}^{2}\approx 1+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$. Using these for equation (4), we obtain
$\mathcal{E}_{\mathrm{I}}$ $=-\mathbf{\hat{B}\cdot}\left(\mathbf{\hat{n}}_{\left[0\right]}\times\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}\right)$
$\approx\frac{4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)+2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}$
$\hskip 10.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (14)
$\mathcal{E}_{\mathrm{II}}$ $=-\mathbf{\hat{B}\cdot}\left(\mathbf{\hat{n}}_{\left[0\right]}\times\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}\right)$
$\approx-\frac{2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right).$ (15)
By means of equations (3), (5) and (8) one can express
$\cos\vartheta$ $\approx\frac{4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (16)
$\sin\vartheta$ $\approx\frac{2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right).$ (17)
Now, using the relations between equations (14), (15) and (16), (17), the birefringent vector can finally be specified from equation (2):
$\hat{\Omega}_{1}$ $\approx-\eta B^{2}\sin^{4}\vartheta+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (18)
$\hat{\Omega}_{2}$ $\approx 2\eta B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (19)
where $\eta\equiv\alpha_{\mathrm{e}}/\left(30\pi B_{\mathrm{c}}^{2}\right)$ and
$B=\frac{B_{\mathrm{max}}r_{*}^{3}\ \left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}{2\left(x^{2}+z^{2}\right)^{3/2}},$ (20)
with $B_{\mathrm{max}}$ being the maximum magnetic field strength at the polar cap [^6] From equation (5) $B_{\mathrm{max}}=\left|B\left(r=r_{*},\theta=\alpha\right)\right|$. and $r_{*}$ being the neutron star radius ($\approx 10^{6}\,\mathrm{cm}$), and $\hat{n}_{y\left[0\right]}$, $\cos\vartheta$ and $\sin\vartheta$ are given by equations (11), (16) and (17), respectively.
To facilitate solving the evolution equations (1) in the next Subsection, we substitute the following identities,
$\cos\left(\theta-\alpha\right)=\frac{\sin\alpha x+\cos\alpha z}{\left(x^{2}+z^{2}\right)^{1/2}},~{}~{}\sin\left(\theta-\alpha\right)=\frac{\cos\alpha x-\sin\alpha z}{\left(x^{2}+z^{2}\right)^{1/2}},~{}~{}\sin\left(\theta-\theta_{\mathrm{o}}\right)=\frac{\cos\theta_{\mathrm{o}}x-\sin\theta_{\mathrm{o}}z}{\left(x^{2}+z^{2}\right)^{1/2}},$ (21)
together with
$x=\hat{n}_{x\left[0\right]}s+r_{\mathrm{o}}\sin\theta_{\mathrm{o}},~{}~{}z=\hat{n}_{z\left[0\right]}s+r_{\mathrm{o}}\cos\theta_{\mathrm{o}}$ (22)
into (16) and (17). Then our solutions for the Stokes vector $\mathbf{S}$ will be parametrised by $s$.
2.2 Solving evolution equations
From equation (1) we write down a system of first-order ordinary differential equations to solve:
$\dot{S}_{1}\left(s\right)$ $=k\hat{\Omega}_{2}\left(s\right)S_{3}\left(s\right),$ (23)
$\dot{S}_{2}\left(s\right)$ $=-k\hat{\Omega}_{1}\left(s\right)S_{3}\left(s\right),$ (24)
$\dot{S}_{3}\left(s\right)$ $=k\left[\hat{\Omega}_{1}\left(s\right)S_{2}\left(s\right)-\hat{\Omega}_{2}\left(s\right)S_{1}\left(s\right)\right],$ (25)
where an over-dot ${\dot{}}$ denotes differentiation with respect to $s$, and $\hat{\Omega}_{1}\left(s\right)$ and $\hat{\Omega}_{2}\left(s\right)$ are given by (18) and (19), respectively. By solving these equations numerically, we find out how the photon polarization evolves through the strong magnetic field in the background of our pulsar emission.
However, in case $\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\ll 1$, one can obtain a solution to equation (1) via perturbation:
$\mathbf{S}=\mathbf{S}_{\left[0\right]}+\delta\mathbf{S}_{\left[1\right]}=\mathbf{S}_{\left[0\right]}+k\int\mathbf{\hat{\Omega}}\times\mathbf{S}_{\left[0\right]}\,\mathrm{d}s,$ (26)
where $\delta\mathbf{S}_{\left[1\right]}$ means the leading order quantum correction to the unperturbed (initial) Stokes vector $\mathbf{S}_{\left[0\right]}$. Here the correction can be treated as the leading order perturbation with $\alpha_{\mathrm{e}}/\left(30\pi\right)\left(B/B_{\mathrm{c}}\right)^{2}\sim 10^{-5}\left(B/B_{\mathrm{c}}\right)^{2}$ being a perturbation parameter. Upon inspection of equations (18) and (19) for equation (26), we can further write down our solution in terms of its components:
$S_{1}$ $\approx S_{1\left[0\right]}+2k\eta S_{3\left[0\right]}\int B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]\mathrm{d}s,$ (27)
$S_{2}$ $\approx S_{2\left[0\right]}+k\eta S_{3\left[0\right]}\int B^{2}\sin^{4}\vartheta\mathrm{d}s,$ (28)
$S_{3}$ $\approx S_{3\left[0\right]}-k\eta\left\{S_{2\left[0\right]}\int B^{2}\sin^{4}\vartheta\mathrm{d}s+2S_{1\left[0\right]}\int B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]\mathrm{d}s\right\}.$ (29)
2.2.1 Examples
We consider X-ray emissions from three neutron stars, (i) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (ii) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (iii) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), and assume for all the emissions, $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$). These stars belong to ‘rotation-powered pulsars’ (RPPs) (Pavlov et al. (2013)). [^7] RPPs refer to neutron stars whose radiation is powered by loss of their rotation energy, via creation and acceleration of $e^{+}e^{-}$ pairs in the strong magnetic field, $B_{\mathrm{max}}\sim 10^{11}-10^{13}\,\mathrm{G}$. The number of detected RPPs are known to be about $\sim 2000$ in radio, $\sim 10$ in optical (including NIR and UV), $\sim 100$ in X-ray and $\sim 100$ in gamma-ray emissions (Pavlov et al. (2013)). In Fig. 2 the three RPPs chosen from the X-ray group are encircled, (i) the one in orange, (ii) another in cyan, (iii) the third in green. Given the X-ray emissions from these, we solve the evolution equations (1) for the following cases.
Figure 2: The population of detected rotation-powered pulsars (RPPs) plotted against their rotation period. RPPs detected in X-rays and gamma-rays are represented by red dots and blue stars, respectively. Three RPPs chosen from the X-ray group, (i) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (ii) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (iii) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), are encircled in orange, cyan and green colours, respectively. (Credit: Pavlov et al. (2013), reproduced with modifications.)
Example (i)
With $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), we obtain numerical solutions through equations (23)-(25) (in solid lines) or perturbative solutions by means of equations (27)-(29) (in dashed lines), as shown in Figs. 3a and 3b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively. The perturbative solutions agree well with numerical ones as $\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\sim\left|k\hat{\Omega}_{1,2\left(\mathrm{min}\right)}s_{\dagger 1,2}\right|\sim 10^{-2}\ll 1$, where $\hat{\Omega}_{1,2\left(\mathrm{min}\right)}\equiv\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)$, evaluated from the birefringent functions given by (18), (19), with $s_{\dagger 1,2}$ being the extremum, i.e., $\mathrm{d}\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)/\mathrm{d}s=0$. On the Poincaré sphere, our solutions are represented by the magenta and light blue loci in Fig. 6a, corresponding to Figs. 3a and 3b, respectively. The loci imply a fraction of an oscillation for the polarization evolution, as is confirmed later by the approximate analytical solutions in Section 2.2.2.
(a) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$
(b) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$
Figure 3: Corresponding to Example (i): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$).
Example (ii)
With $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), we obtain numerical solutions through equations (23)-(25), as shown in Figs. 4a and 4b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively. On the Poincaré sphere, our solutions are represented by the magenta and light blue loci in Fig. 6b, corresponding to Figs. 4a and 4b, respectively. The loci imply about half an oscillation for the polarization evolution, as is confirmed later by the approximate analytical solutions in Section 2.2.2.
(a) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$
(b) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$
Figure 4: Corresponding to Example (ii): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$).
Example (iii)
With $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), we obtain numerical solutions through equations (23)-(25), as shown in Figs. 5a and 5b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively. On the Poincaré sphere, our solutions are represented by the magenta and light blue loci in Fig. 6c, corresponding to Figs. 5a and 5b, respectively. The loci imply multiple oscillations for the polarization evolution, as is confirmed later by the approximate analytical solutions in Section 2.2.2.
(a) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$
(b) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$
Figure 5: Corresponding to Example (iii): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$).
(a) Corresponding to Example (i):
a fraction of an oscillation
(b) Corresponding to Example (ii):
about half an oscillation
(c) Corresponding to Example (iii):
multiple oscillations
Figure 6: Representations of the Stokes vectors on the Poincaré sphere. The loci imply patterns of the polarization evolution in terms of oscillation.
2.2.2 Approximate analytical solutions
Plotting the birefringent functions $\hat{\Omega}_{1}\left(s\right)$ and $\hat{\Omega}_{2}\left(s\right)$, as given by (18) and (19), respectively, one can observe that they feature distinctive patterns; they can be well approximated by some analytic models, whose curves resemble the original plots. In Fig. 7 are plotted the birefringent functions for the three cases: (a) $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (b) $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (c) $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$) with solid lines (see Figs. 7a, 7b and 7c, respectively), where they have been evaluated with the same initial condition as assumed in Section 2.2.1.
In correspondence with the actual birefringent functions above, the following analytical models are also plotted with dashed lines in Fig. 7:
$\hat{\Omega}_{1,2}\left(s\right)\approx-a_{1,2}s^{\frac{p+1}{p}}e^{-bs}~{}~{}\text{for $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$},$ (30)
where $a_{1,2}>0$, $b>0$ and $p>0$ are free parameters; with suitable values chosen for these, our model functions can give rise to solutions of equations (23)-(25) that match fairly well the numerical results obtained in Section 2.2.1. Here we express
$a_{1,2}=-\hat{\Omega}_{1,2\left(\mathrm{min}\right)}\left[\frac{e}{\left(1-\chi\right)s_{\dagger 1}+\chi s_{\dagger 2}}\right]^{\frac{p+1}{p}},~{}~{}b=\frac{p+1}{p\left[\left(1-\chi\right)s_{\dagger 1}+\chi s_{\dagger 2}\right]},$ (31)
where $\hat{\Omega}_{1,2\left(\mathrm{min}\right)}=\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)$, evaluated from (18), (19), with $s_{\dagger 1,2}$ denoting the extremum. We have set $p=100$ (a sufficiently large number) for all three cases, and $\chi=0.999$ for (a) and (b), and $\chi=0.056$ for (c) in Fig. 7.888In particular, the values for $\chi$ have been chosen such that our solutions converge to the asymptotic limits that match well the numerical results given by Figs. 3b, 4b and 5b in Section 2.2.1, as $s$ tends to $\infty$.
(a) For $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$)
(b) For $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$)
(c) For $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$)
Figure 7: Plots of the birefringent functions $\hat{\Omega}_{1}\left(s\right)$ and $\hat{\Omega}_{2}\left(s\right)$ and their approximate analytic models.
Plugging equation (30) into equations (23)-(25), we obtain analytical solutions as follows (for a complete derivation, see Appendix B):
$S_{1}\left(s\right)$ $\approx\frac{a_{2}S_{\mathrm{o}}}{\sqrt{a_{1}^{2}+a_{2}^{2}}}\cos\left(\Psi\left(s;p\right)+\delta\right)+\frac{a_{1}C}{\sqrt{a_{1}^{2}+a_{2}^{2}}},$ (32)
$S_{2}\left(s\right)$ $\approx-\frac{a_{1}S_{\mathrm{o}}}{\sqrt{a_{1}^{2}+a_{2}^{2}}}\cos\left(\Psi\left(s;p\right)+\delta\right)+\frac{a_{2}C}{\sqrt{a_{1}^{2}+a_{2}^{2}}},$ (33)
$S_{3}\left(s\right)$ $\approx S_{\mathrm{o}}\sin\left(\Psi\left(s;p\right)+\delta\right),$ (34)
where
$\Psi\left(s;p\right)\equiv k\sqrt{a_{1}^{2}+a_{2}^{2}}b^{-\frac{4p+1}{2p}}s^{\frac{1}{2p}}e^{-\frac{1}{2}bs}\left[M_{\frac{1}{2p},\frac{p+1}{2p}}\left(bs\right)-\left(bs\right)^{\frac{2p+1}{2p}}e^{-\frac{1}{2}bs}\right],$ (35)
and $M_{\kappa,\mu}\left(z\right)$ denotes a Whittaker function of the first kind, [^9] The expression inside the square brackets in equation (35) has been reduced from its original form as given by equation (60) in Appendix B, using the identity $M_{\left(2p+1\right)/\left(2p\right),\left(p+1\right)/\left(2p\right)}\left(bs\right)=\left(bs\right)^{\left(2p+1\right)/\left(2p\right)}e^{-bs/2}M\left(0,\left(2p+1\right)/p,bs\right)$, with the Kummer function $M\left(0,\left(2p+1\right)/p,bs\right)=1$ as a special case (NIST (2024)). and $a_{1,2}$ and $b$ are given by (31). Here we determine $S_{\mathrm{o}}$, $C$ and $\delta$ by matching the initial value of the Stokes vector $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)$ with equations (32)-(34) evaluated at $s=0$.
In Fig. 8 we plot the above solutions for the following cases, assuming $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ for $\hat{\Omega}_{1,2\left(\mathrm{min}\right)}$ (evaluated through (18), (19)) and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$) for the X-ray emissions, with the initial Stokes vector $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$:
Case (a) (see Fig. 8a);
for $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), with the parameters
$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-2.5511\times 10^{-16}$, $\hat{\Omega}_{2\left(\mathrm{min}\right)}\approx-2.2054\times 10^{-16}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $\chi=0.999$, $S_{\mathrm{o}}\approx 0.7960$, $C\approx 0.6052$ and $\delta\approx 0.8537$,
Case (b) (see Fig. 8b);
for $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), with the parameters
$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-8.0001\times 10^{-15}$, $\hat{\Omega}_{2\left(\mathrm{min}\right)}\approx-3.9240\times 10^{-16}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $\chi=0.999$, $S_{\mathrm{o}}\approx 0.6013$, $C\approx 0.7990$ and $\delta\approx 1.5056$,
Case (c) (see Fig. 8c);
for $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), with the parameters
$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-6.3776\times 10^{-13}$, $\hat{\Omega}_{2\left(\mathrm{min}\right)}\approx-2.7567\times 10^{-14}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $\chi=0.056$, $S_{\mathrm{o}}\approx 0.6010$, $C\approx 0.7993$ and $\delta\approx 1.5133$.
These plots compare with Figs. 3b (for Example (i)), 4b (for Example (ii)) and 5b (for Example (iii)) in Section 2.2.1, respectively.
(a) For $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$)
(b) For $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$)
(c) For $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$)
Figure 8: Plots of the analytical solutions $S_{1}\left(s\right)$, $S_{2}\left(s\right)$ and $S_{3}\left(s\right)$, given the initial Stokes vector $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$.
The analytical solutions (32)-(34) provide a useful tool for understanding the different patterns of polarization evolution for the three cases above, as given by Figs. 8a, 8b and 8c. Inspecting numerically the functional argument $\Psi\left(s;p\right)$ given by (35), one can approximate it to a simpler form with the help of (30) and (31):
For $0\leq s\lesssim 2.5\times 10^{6}\,\mathrm{cm}$,
$\Psi\left(s;p=100\right)=k\sqrt{a_{1}^{2}+a_{2}^{2}}b^{-\frac{401}{200}}s^{\frac{1}{200}}e^{-\frac{1}{2}bs}\left[M_{\frac{1}{200},\frac{101}{200}}\left(bs\right)-\left(bs\right)^{\frac{201}{200}}e^{-\frac{1}{2}bs}\right]\approx 0.24k\sqrt{\hat{\Omega}_{1\left(\mathrm{min}\right)}^{2}+\hat{\Omega}_{2\left(\mathrm{min}\right)}^{2}}\pi s.$ (36)
Using this, we can estimate how much the oscillations for the three cases have progressed, for example, during $0\leq s\lesssim 1.2\times 10^{6}\,\mathrm{cm}$:
$\Psi\left(s\approx 1.2\times 10^{6}\,\mathrm{cm};p=100\right)$$\approx\left\{\begin{array}[]{ll}0.02\pi\hskip 9.0pt\text{(a fraction of an oscillation)}&\hskip 12.0pt\text{for Case (a)},\\ 0.5\pi\hskip 13.0pt\text{(about half an oscillation)}&\hskip 12.0pt\text{for Case (b)},\\ 39\pi\hskip 15.0pt\text{(multiple oscillations)}&\hskip 12.0pt\text{for Case (c)},\end{array}\right.$ (40)
which can be checked by comparison with Figs. 8a, 8b and 8c, respectively.
Recalling Example (i) from Section 2.2.1, one finds that the condition for perturbation can be equivalently expressed using (36):
[…]
Extending this argument, similarly to (40) above we may state the following in reference to the patterns of polarization evolution:
$\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\left\{\begin{array}[]{ll}\ll 1\hskip 9.0pt\text{(fractionally oscillatory - monotonic)}&\hskip 12.0pt\text{for Example (i)},\\ \sim\hskip 2.0pt1\hskip 9.0pt\text{(half oscillatory)}&\hskip 12.0pt\text{for Example (ii)},\\ \gg 1\hskip 9.0pt\text{(highly oscillatory)}&\hskip 12.0pt\text{for Example (iii)}.\end{array}\right.$ (45)
3 Conclusions and discussion
We have investigated the evolution of polarization states under the quantum refraction effects combined with the frequency dependence of pulsar emission. To this end, we have solved a system of evolution equations of the Stokes vector given by (1) (or by (23)-(25)) for three examples of RPPs at a fixed frequency for specific emissions (e.g., X-rays as in Sections 2.2.1 and 2.2.2). Our main results are presented in Figs. 3-6, from numerical solutions and in part from perturbative solutions. Also, we have approximated the birefringent vector by some models as in Fig. 7 to solve the evolution equations analytically, and obtained the results as presented in Fig. 8, from approximate analytical solutions. It is noteworthy that at a fixed frequency of emission the evolution of the Stokes vector largely exhibits three different patterns, depending on the magnitudes of the birefringent vector, in which the magnetic field strength is a dominant factor: (i) fractionally oscillatory - adiabatic/monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviours. These features are shown by the numerical solutions in Figs. 3-6, and also confirmed by the approximate analytical solutions in Fig. 8.
This study is centred on solving the evolution equations for polarization states (1), wherein the birefringent vector that contains all information about the quantum refraction effects, coupled to the frequency of pulsar emission, acts on the Stokes vector; the evolution results from the combination of the quantum refraction effects and the frequency dependence of the emission. This is a major difference from our previous work (Kim et al. (2024)), wherein the same effects are independent of the emission frequency; the work solely focuses on the quantum refraction effects on the propagation and polarization of a photon from pulsar emission, with no reference to other properties of the emission, such as the frequency. In this regard, it is worthwhile to draw comparison between the two quantities, the polarization vector and the Stokes vector, both of which are used to describe polarization states. The polarization vector is defined directly from the radiative electric field vector (i.e. the unit electric field vector), and it is parallel-transported along the the propagation vector; usually, we consider such two vectors orthogonal to each other and to the propagation vector to define an orthonormal basis consisting of the three vectors. In contrast, the Stokes vector is defined from Stokes parameters which are built out of the radiative electric field vector (Rybicki & Lightman (1991)). The representation of the Stokes vector is abstract in the sense that it is a vector defined on the Poincaré sphere. The Stokes vector is not parallel-transported along the the propagation vector, but can still be defined along the propagation vector as the two polarization vectors move along it; hence, it can be parametrised by $s$ to represent polarization states along the photon trajectory. However, the Stokes vector has a crucial advantage over the polarization vector in representing polarization states in some astrophysical studies like this: it can be directly estimated from polarimetric measurements and accommodate depolarization effects due to incomplete coherence and random processes during the photon propagation (Scully & Zubairy (1997)).
With regard to the adiabatic evolution condition as mentioned in Heyl & Shaviv (2000), Heyl et al. (2003), and Heyl & Caiazzo (2018), we have carefully examined our results presented in Figs. 3-5 to see what interpretations the condition leads to. Solving the condition $\left|k\mathbf{\hat{\Omega}}\left(\mathrm{d}\ln|k\mathbf{\hat{\Omega}}|/\mathrm{d}s\right)^{-1}\right|\gtrsim 0.05$ for $s$ (note that our $k\mathbf{\hat{\Omega}}$ is equivalent to the birefringent vector as defined in the references above and that we have set the condition value to $0.05$ rather than $0.5$ as in the references) yields $s_{\mathrm{PL}1}\lesssim s\lesssim s_{\mathrm{PL}2}$, where $s_{\mathrm{PL}1[2]}$ refers to the lower [upper] bound for the ‘polarization limiting’ distance as measured from the emission point. Using this, we have checked out the following: (1) for $6.2\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 1.8\times 10^{6}\,\mathrm{cm}$ in Fig. 3, (2) for $3.1\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 6.2\times 10^{6}\,\mathrm{cm}$ in Fig. 4, (3) for $1.1\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 1.9\times 10^{7}\,\mathrm{cm}$ in Fig. 5, our Stokes vector evolves evidently; otherwise, it freezes.
Effects of gravitation have not been considered in this study. However, close to the neutron star, where gravitation due to the neutron star mass may not be negligible, its effects must be taken into account in our analysis. Then, basically, the following shall be redefined in curved spacetime: (1) the QED one-loop effective Lagrangian, (2) the refractive index for the photon propagation, (3) the magnetic field geometry in the magnetosphere, (4) the radiative electric field due to a charge moving along a magnetic field line, (5) the photon trajectory. All these have not been rigorously dealt with in previous studies. In this regard, inclusion of the gravitational effects will involve non-trivial and immense analyses, and therefore shall be conducted for a long-term plan in our future studies.
Acknowledgements
D.-H.K was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1I1A1A01054781). C.M.K. was supported by Ultrashort Quantum Beam Facility operation program (140011) through APRI, GIST and GIST Research Institute (GRI) grant funded by GIST. S.P.K. was also in part supported by National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2019R1I1A3A01063183).
Data Availability
The inclusion of a Data Availability Statement is a requirement for articles published in MNRAS. Data Availability Statements provide a standardised format for readers to understand the availability of data underlying the research results described in the article. The statement may refer to original data generated in the course of the study or to third-party data analysed in the article. The statement should describe and provide means of access, where possible, by linking to the data or providing the required accession numbers for the relevant databases or DOIs.
Appendix A The classical Stokes vector
Consider a particle with a charge $q$ moving along a curved trajectory (a magnetic field line). Then the curvature radiation due to this can be expressed by the electric field:
$\mathbf{E}\left(t\right)=\frac{q}{c\left|\mathbf{r}-\boldsymbol{\xi}\left(t_{\mathrm{ret}}\right)\right|}\frac{\mathbf{n\times}\left[\left(\mathbf{n-}\frac{\dot{\boldsymbol{\xi}}\left(t_{\mathrm{ret}}\right)}{c}\right)\times\frac{\ddot{\boldsymbol{\xi}}\left(t_{\mathrm{ret}}\right)}{c}\right]}{\left(1-\frac{\dot{\boldsymbol{\xi}}\left(t_{\mathrm{ret}}\right)}{c}\cdot\mathbf{n}\right)^{3}},$ (46)
where $t_{\mathrm{ret}}\equiv t-r/c$ is the retarded time, $\boldsymbol{\xi}$ represents the particle’s trajectory, $\mathbf{n}$ is the propagation direction of the radiation, and an over-dot ${\dot{}}$ denotes differentiation with respect to $t$. In a suitably chosen Cartesian frame, by setting $\boldsymbol{\xi}\left(t_{\mathrm{ret}}\right)=\rho\left(\sin\left(\beta ct_{\mathrm{ret}}/\rho\right),0,\cos\left(\beta ct_{\mathrm{ret}}/\rho\right)\right)$, with $\rho$ being the radius of curvature of the particle’s trajectory, and $\mathbf{n}=\left(\cos\varphi,\sin\varphi,0\right)$, with $\varphi$ being the angle measured from the $x$-axis to the plane of the particle’s motion, we can construct a simple toy model for pulse profiles of pulsar curvature emission as described below (Kim & Trippe (2021)).
One can express Stokes parameters out of the radiation field (46), which describe polarization properties of the curvature radiation (Gil & Snakowski (1990)):
$I$ $=\tilde{E}_{\parallel}^{\ast}\tilde{E}_{\parallel}+\tilde{E}_{\perp}^{\ast}\tilde{E}_{\perp}$
$=\mathcal{E}_{\mathrm{o}}^{2}\omega^{2}\left[\left(\delta^{2}+\varphi^{2}\right)^{2}\mathrm{K}_{2/3}^{2}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)+\,\varphi^{2}\left(\delta^{2}+\varphi^{2}\right)\mathrm{K}_{1/3}^{2}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)\right],$ (47)
$Q$ $=\tilde{E}_{\parallel}^{\ast}\tilde{E}_{\parallel}-\tilde{E}_{\perp}^{\ast}\tilde{E}_{\perp}$
$=\mathcal{E}_{\mathrm{o}}^{2}\omega^{2}\left[\left(\delta^{2}+\varphi^{2}\right)^{2}\mathrm{K}_{2/3}^{2}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)-\,\varphi^{2}\left(\delta^{2}+\varphi^{2}\right)\mathrm{K}_{1/3}^{2}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)\right],$ (48)
$U$ $=\tilde{E}_{\parallel}^{\ast}\tilde{E}_{\perp}+\tilde{E}_{\parallel}\tilde{E}_{\perp}^{\ast}=0,$ (49)
$V$ $=-\mathrm{i}\left(\tilde{E}_{\parallel}^{\ast}\tilde{E}_{\perp}-\tilde{E}_{\parallel}\tilde{E}_{\perp}^{\ast}\right)$
$=-2\mathcal{E}_{\mathrm{o}}^{2}\omega^{2}\varphi\left(\delta^{2}+\varphi^{2}\right)^{3/2}\mathrm{K}_{2/3}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)\mathrm{K}_{1/3}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right),$ (50)
where $\tilde{E}_{\parallel}$ and $\tilde{E}_{\perp}$ denote the components of the Fourier transform $\mathbf{\tilde{E}}\left(\omega\right)=\tilde{E}_{\parallel}\left(\omega\right)\mathbf{e}_{z}+\tilde{E}_{\perp}\left(\omega\right)\,\mathbf{e}_{y}=\int_{-\infty}^{\infty}\mathbf{E}\left(t\right)\exp\left(\mathrm{i}\omega t\right)\,\mathrm{d}t$, expressed in the Cartesian frame, and means the complex conjugate, and $\mathcal{E}_{\mathrm{o}}=q\beta/\left(2\sqrt{3}\pi^{2}r\rho\right)$, and $\delta\equiv\gamma^{-1}=\left(1-\beta^{2}\right)^{1/2}\ll 1$ is the half-angle of the beam emission, and $\mathrm{K}_{1/3}$ and $\mathrm{K}_{2/3}$ denote the modified Bessel functions of the second kind. With regard to the polarization state of the curvature radiation, $I$ is a measure of the total intensity, $Q$ and $U$ jointly describe the linear polarization, and $V$ describes the circular polarization. These parameters can be plotted as functions of the phase angle $\varphi$, where $\varphi\leq\delta\ll 5^{\circ}$ usually, to simulate the pulse profiles of pulsar emission theoretically.
Out of the Stokes parameters, one can define the Stokes vector $\mathbf{S}=\left(S_{1},S_{2},S_{3}\right)\allowbreak\equiv\left(Q/I,U/I,V/I\right)$ and express it using (47)-(50):
$S_{1}$ $=\frac{\left(\delta^{2}+\varphi^{2}\right)\mathrm{K}_{2/3}^{2}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)-\,\varphi^{2}\mathrm{K}_{1/3}^{2}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)}{\left(\delta^{2}+\varphi^{2}\right)\mathrm{K}_{2/3}^{2}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)+\,\varphi^{2}\mathrm{K}_{1/3}^{2}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)},$ (51)
$S_{2}$ $=0,$ (52)
$S_{3}$ $=-\frac{2\varphi\left(\delta^{2}+\varphi^{2}\right)^{1/2}\mathrm{K}_{2/3}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)\mathrm{K}_{1/3}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)}{\left(\delta^{2}+\varphi^{2}\right)\mathrm{K}_{2/3}^{2}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)+\,\varphi^{2}\mathrm{K}_{1/3}^{2}\left(\frac{\rho\omega}{3\beta c}\left(\delta^{2}+\varphi^{2}\right)^{3/2}\right)}.$ (53)
In Fig. 9 is plotted the classical Stokes vector against the phase angle $\varphi$, where we have set, for example, $\delta\approx 10^{-8}$, $\rho\approx 10^{8}\,\mathrm{cm}$, $\beta\approx 1$ and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ to model pulse profiles of X-ray pulsar emission. Here the initial values for the Stokes vector, $\mathbf{S}\left(\varphi=0\right)=\left(1,0,0\right)$ and $\mathbf{S}\left(\varphi\approx-1.16\times 10^{-6}\,\mathrm{rad}\right)=\left(0.8,0,0.6\right)$, as in the examples given in Section 2.2.1, are marked by solid circles and solid boxes, respectively.
Figure 9: The classical Stokes vector $\mathbf{S}\left(\varphi\right)=\left(S_{1}\left(\varphi\right),S_{2}\left(\varphi\right),S_{3}\left(\varphi\right)\right)$ plotted against the phase angle $\varphi$. Its initial values $\mathbf{S}\left(\varphi=0\right)=\left(1,0,0\right)$ and $\mathbf{S}\left(\varphi\approx-1.16\times 10^{-6}\,\mathrm{rad}\right)=\left(0.8,0,0.6\right)$ are marked by solid circles and solid boxes, respectively.
Appendix B Approximate analytical solutions to evolution equations
Substituting equation (30) into equations (23)-(25), the evolution equations can be reduced as follows:
For $0\leq s\leq 20r_{*}$,
$\dot{S}_{1}\left(s\right)\approx-ka_{2}s^{\frac{p+1}{p}}e^{-bs}S_{3}\left(s\right),$ (54)
$\dot{S}_{2}\left(s\right)\approx ka_{1}s^{\frac{p+1}{p}}e^{-bs}S_{3}\left(s\right),$ (55)
$\ddot{S}_{3}\left(s\right)-\left(\frac{p+1}{ps}-b\right)\dot{S}_{3}\left(s\right)+k^{2}\left(a_{1}^{2}+a_{2}^{2}\right)\left(s^{\frac{p+1}{p}}e^{-bs}\right)^{2}S_{3}\left(s\right)\approx 0.$ (56)
First, we solve equation (56) for $S_{3}\left(s\right)$, and then using this solution, obtain $S_{1}\left(s\right)$ and $S_{2}\left(s\right)$, by integrating equations (54) and (55), respectively:
$S_{3}\left(s\right)$ $\approx S_{\mathrm{o}}\sin\left(\Psi\left(s;p\right)+\delta\right),$ (57)
$S_{1}\left(s\right)$ $\approx\frac{a_{2}S_{\mathrm{o}}}{\sqrt{a_{1}^{2}+a_{2}^{2}}}\cos\left(\Psi\left(s;p\right)+\delta\right)+S_{1\mathrm{o}},$ (58)
$S_{2}\left(s\right)$ $\approx-\frac{a_{1}S_{\mathrm{o}}}{\sqrt{a_{1}^{2}+a_{2}^{2}}}\cos\left(\Psi\left(s;p\right)+\delta\right)+S_{2\mathrm{o}},$ (59)
where
$\Psi\left(s;p\right)\equiv k\sqrt{a_{1}^{2}+a_{2}^{2}}b^{-\frac{4p+1}{2p}}s^{\frac{1}{2p}}e^{-\frac{1}{2}bs}\left[M_{\frac{1}{2p},\frac{p+1}{2p}}\left(bs\right)-M_{\frac{2p+1}{2p},\frac{p+1}{2p}}\left(bs\right)\right],$ (60)
and $M_{\kappa,\mu}\left(z\right)$ denotes a Whittaker function of the first kind. Here employing the identity $S_{1}^{2}\left(s\right)+S_{2}^{2}\left(s\right)+S_{3}^{2}\left(s\right)=1$ (conservation of the degree of polarization), one can specify $S_{1\mathrm{o}}$ and $S_{2\mathrm{o}}$ in terms of $a_{1}$, $a_{2}$ and a constant $C$, and establish a relation between $S_{\mathrm{o}}$ and $C$:
$S_{1\mathrm{o}}=\frac{Ca_{1}}{\sqrt{a_{1}^{2}+a_{2}^{2}}},~{}~{}S_{2\mathrm{o}}=\frac{Ca_{2}}{\sqrt{a_{1}^{2}+a_{2}^{2}}},~{}~{}S_{\mathrm{o}}^{2}+C^{2}=1.$ (61)
Then $S_{\mathrm{o}}$, $C$ and $\delta$ are determined by matching the initial value of the Stokes vector $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)$ with equations (57)-(59) evaluated at $s=0$.

### v3

@@
-The quantum-refractive evolution of polarization states in pulsar emission
+
 Abstract
 Highly magnetized neutron stars have quantum refraction effects on pulsar emission due to the non-linearity of the quantum electrodynamics (QED) action.
-In this paper, we investigate the evolution of the polarization states under the quantum refraction effects combined with the frequency dependence of pulsar emission; we solve a system of evolution equations of the Stokes vector, where the birefringent vector, in which such effects are encoded, acts on the Stokes vector.
-At a fixed frequency of emission, depending on the magnitude of the birefringent vector, dominated mostly by the magnetic field strength, the evolution of the Stokes vector largely exhibits three different patterns: (i) monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviours.
+In this paper, we investigate the evolution of the polarization states of pulsar emission under the quantum refraction effects, combined with the dependence on the emission frequency, for dipole and quadrudipole pulsar models; we solve a system of evolution equations of the Stokes vector, where the birefringent vector, in which such effects are encoded, acts on the Stokes vector.
+At a fixed emission frequency, depending on the magnitude of the birefringent vector, dominated mostly by the magnetic field strength, the evolution of the Stokes vector largely exhibits three different patterns: (i) monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviors.
 These features are understood and confirmed by means of approximate analytical solutions to the evolution equations.
+Also, the evolution patterns are shown to differ between dipole and quadrudipole pulsar models, depending on the magnetic field strength.
 keywords:
-pulsars – magnetic fields – photon emission – non-linear electrodynamics – quantum refraction – polarization evolution
-pubyear: 2024pagerange: The quantum-refractive evolution of polarization states in pulsar emissionB
+pulsars, magnetic fields, photon emission, photon polarization, nonlinear electrodynamics, nonlinear vacuum, vacuum birefringence, quantum refraction, polarization evolution
 1 Introduction
-Highly magnetized neutron stars are interesting astrophysical objects that will provide an astrophysical arena for probing fundamental physics (Raffelt (1996)), in particular, strong field QED (Meszaros (1992); Harding & Lai (2006)).
-The magnetic fields of neutron stars of $10^{12}-10^{15}\,\mathrm{G}$ are strong enough to give rise to QED effects, such as vacuum polarization and pair production due to the induced electric fields, which have been proposed as an astrophysical laboratory to test strong field QED (see Kim & Kim (2021) and references therein).
-A great advantage of neutron stars as a laboratory is virtually infinitely long duration and range of the electromagnetic fields (in the scale of Compton time and length), strongly contrasting with the short, pulsed fields achieved with ultra-intense lasers, as well as the strength of the fields (Yoon et al.
-(2021)).
-Strong electromagnetic fields polarize the vacuum, which leads to the Heisenberg-Euler-Schwinger (HES) one-loop effective action (Heisenberg & Euler (1936); Schwinger (1951)).
-The effective action describes non-linear electrodynamics, which has an imaginary part from the vacuum instability due to production of charged particle pairs via electric fields.
-The HES action leads to the vacuum multiple refraction indices, which is generally referred to as the quantum refraction effect.
-A test of strong-field QED is one of the most important problems in physics, and there have been laboratory experiment proposals for it, e.g., PVLAS (Polarizzazione del Vuoto con LAser) project (Ejlli et al.
-(2020)).
-On the other hand, highly magnetized neutron stars provide near-critical fields, and these can be used as various celestial laboratories to test strong field QED (Ruffini et al.
-(2010); Fedotov et al.
-(2023); Hattori et al.
-(2023))).
-Also, there have been space missions proposed to observe the strong-field QED effects, e.g., the enhanced X-ray Timing and Polarimetry (eXTP) (Santangelo et al.
-(2019)) and the Compton Telescope project (Wadiasingh et al.
-(2019)).
-The HES action is well approximated by the post-Maxwellian action, even up to the strength one order lower than the critical magnetic field $B_{\mathrm{c}}=m_{\mathrm{e}}^{2}c^{3}/e\hbar=4.4\times 10^{13}\,\mathrm{G}$, which keeps up to the quadratic terms of the Maxwell scalar and pseudo-scalar.
-Therefore, the post-Maxwellian action exhibits non-linear characteristics of vacuum polarization, such as quantum refraction.
-Previously, we have studied the quantum refraction effects on the polarization vectors in the dipole magnetic field background of a pulsar model (Kim et al.
-(2024)).
-The study is non-trivial in comparison with other similar studies where the background magnetic field is assumed to be uniform, in that we have to deal with dipole magnetic field, whose strength and direction vary over space.
-In this paper, we study the evolution of the polarization states under the quantum refraction effects combined with the frequency dependence of pulsar emission.
-Our analysis is based on the evolution equations of the Stokes vector, where such effects are encoded into the birefringent vector that acts on the Stokes vector.
+Strong fields may open a window for testing fundamental physics.
+Even before quantum electrodynamics (QED) was fully developed and precisely tested in the weak field regime, Heisenberg and Euler showed that a strong electromagnetic field can polarize the Dirac vacuum Heisenberg1936Folgerungen .
+Schwinger introduced the proper-time integral method to obtain the one-loop effective QED action of the vacuum under a uniform electromagnetic field Schwinger1951Gauge .
+The so-called Heisenberg-Euler-Schwinger (HES) action provides an effective theory of electrodynamics in strong fields, in which the linear Maxwell vacuum is turned into a dielectric medium with electric, magnetic, and magneto-electric responses.
+Consequently, a photon propagating in a region of strong electromagnetic fields can experience vacuum birefringence, i.e., a quantum refraction effect Meszaros1992High ; Kim2023Vacuum .
+Furthermore, when the electric field is sufficiently strong to be comparable to the critical electric field strength ($E_{\rm{c}}=m_{\mathrm{e}}^{2}c^{3}/\left(e\hbar\right)\approx 1.3\times 10^{16}\,\rm{V/cm}$), electron-positron pairs can be created spontaneously out of the vacuum, which is called Sauter-Schwinger pair production Ruffini_2010 ; Fedotov:2022ely ; Hattori:2023egw .
+Observing these effects will validate the quantum vacuum model in the strong-field regime; however, vacuum birefringence is practically far more likely to be implemented or observed than Sauter-Schwinger pair production.
+Although the relevant field strength is too high to attain by terrestrial means, an experiment and several proposals for testing strong-field QED have been reported.
+In the PVLAS (Polarizzazione del Vuoto con Laser, i.e., polarization of vacuum with laser) project, permanent, superconducting magnets have been used with laser as a probe, and a limit on vacuum birefringence has been reported for a field strength of $2.5\times 10^{4}\,\mathrm{G}$ Ejlli2020PVLASa .
+Recently, several proposals have appeared, in which ultra-intense laser fields are used with X-rays as a probe Karbstein2021Vacuum ; Shen2018Exploring ; Yu2023X ; the current ultra-intense laser can provide a magnetic field strength of $10^{10}\,\rm{G}$ Danson2019Petawatt , which corresponds to millisecond pulsars, but not to young pulsars.
+Although the fields from such lasers are still weaker than the critical magnetic field strength, $B_{\mathrm{c}}=m_{\mathrm{e}}^{2}c^{3}/e\hbar\approx 4.414\times 10^{13}\,\mathrm{G}$ by three orders, these proposals are promising for the observation of vacuum birefringence in the relatively weak-field regime.
+Thus, they are seriously considered to be conducted at upcoming ultra-intense laser facilities Danson2019Petawatt .
+However, the observation of vacuum birefringence in the strong-field regime requires a field strength comparable to the critical value.
+It has been predicted that such extreme fields are available from astrophysical compact objects.
+For instance, highly magnetized neutron stars have magnetospheres whose field strength approaches up to $\sim 2\times 10^{15}\,\rm{G}$ (about 50 times as high as the critical field strength) Meszaros1992High ; Harding2006Physics .
+In this regard, several space telescope missions are being conducted or proposed to observe the X-rays from neutron stars for vacuum birefringence: the Imaging X-ray Polarimetry Explorer (IXPE) IXPE , X-ray Polarimeter Satellite (XPoSat) XPoSat , the enhanced X-ray Timing and Polarimetry (eXTP) Santangelo_2019 and the Compton Telescope project Wadiasingh2019Magnetars .
+The X-rays from a neutron star contain information about vacuum birefringence in its magnetosphere, and the birefringence effect accumulates over the magnetospheric size.
+Such accumulation is a great advantage compared to terrestrial laser experiments, not to mention the available field strength.
+This way, astrophysical compact objects can be used as a laboratory to test fundamental physics in the strong-field regime Raffelt1996Stars ; Meszaros1992High ; Kim2024Magnetars .
+The HES action is well approximated by the post-Maxwellian action, even up to the strength one order lower than the critical magnetic field $B_{\mathrm{c}}$, which keeps up to the quadratic terms of the Maxwell scalar and pseudo-scalar.
+Therefore, the post-Maxwellian action exhibits non-linear characteristics of vacuum polarization, such as quantum refraction Adler1971Photon ; Ni2013Foundations ; Denisov2016Pulsar .
+Previously, we have studied the quantum refraction effects on the propagation of a probe photon in the magnetic dipole field background of a pulsar model 10.1093/mnras/stae1304 .
+The study is non-trivial in comparison with other similar studies wherein the background magnetic field is assumed to be uniform, in that we have to deal with a dipole magnetic field, the strength and direction of which vary over space.
+In this work, we investigate the evolution of the polarization states of pulsar emission under the quantum refraction effects, combined with the dependence on the emission frequency, for dipole and quadrudipole (for the first time, to our knowledge) pulsar models; with growing theoretical and observational concerns for beyond-dipole effects, this study extends the scope of previous works by taking into account the multipolar magnetic field structure, the importance of which has been illuminated in different contexts of pulsar astronomy by a number of studies (see Gralla_2016 ; 10.1093/mnras/stz2524 ; Kazmierczak2019 ; Kalapotharakos_2021 , Petri2016Theory and references therein).
+To this end, we employ the evolution equations of the Stokes vector, where such effects are encoded into the birefringent vector that acts on the Stokes vector.
 The Stokes vector has a crucial advantage over the polarization vector in representing polarization states: it can be directly determined from experimentally measurable quantities and accommodate depolarization effects due to incomplete coherence and random processes during the photon propagation.
 Solutions of the evolution equations describe how the polarization states change along the photon propagation path from the emission point towards an observer.
-It turns out that the evolution of the Stokes vector, at a fixed frequency of emission, largely exhibits three different patterns, depending on the magnitudes of the birefringent vector, dominated mostly by the magnetic field strength: (i) fractionally oscillatory - monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviours, which are found by numerical solutions and also confirmed by approximate analytical solutions.
+It turns out that the evolution of the Stokes vector, at a fixed frequency of emission, largely exhibits three different patterns, depending on the magnitudes of the birefringent vector, dominated mostly by the magnetic field strength: (i) fractionally oscillatory - monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviors, which are found by numerical solutions and also confirmed by approximate analytical solutions.
 These are novel features rarely illuminated in previous studies on the same topic.
+In addition, it is investigated how the aforementioned features regarding the evolution of the Stokes vector change as we replace a dipole field with a quadrudipole field to modify the pulsar magnetic field structure.
+Throughout our analysis, X-ray emission from pulsars, with frequency $\sim 10^{18}\,\mathrm{Hz}$, is considered; in this regime, the vacuum contribution to the birefringence dominates that of the plasma Heyl2000 ; Wang2007Wave .
+Also, our analysis is sufficiently rigorous in solving the evolution equations of the Stokes vector, in that we feed into the equations the precise information of photon propagation under the pulsar rotation effect, through the magnetic field geometries of oblique dipole and quadrudipole rotators, with all the quantities involved fully affine-parameterized; then, the equations are solved solely in terms of an affine parameter.
 The paper is organized as follows.
-In Section 2.1, we introduce a system of evolution equations of the Stokes vector and apply this formalism to our pulsar emission model for an oblique rotator.
-In Section 2.2, the evolution equations are solved for some known rotation-powered pulsars (RPPs) in three ways: fully numerically, via perturbation analysis, and using an analytical approximation.
+In Sect.
+2.1, we introduce a system of evolution equations of the Stokes vector and apply this formalism to our pulsar emission model for an oblique dipole rotator.
+In Sect.
+2.2, the evolution equations are solved for some known rotation-powered pulsars (RPPs) in three ways: fully numerically, via perturbation analysis, and using an analytical approximation.
 Also, we discuss the evolution patterns of the Stokes vectors resulting from the solutions.
-Then finally, we conclude the paper with discussions on other similar studies and follow-up studies.
-2 Evolution of Stokes vector in strong magnetic field
+In Sect.
+3.1, we consider a magnetic quadrudipole model for pulsar emission and look into the evolution equations under this model.
+In Sect.
+3.2, we solve the evolution equations for the same RPPs fully numerically, and compare the results with those for the dipole case in Sect.
+2.2.
+Then finally, we conclude the paper with discussions on other similar studies and future follow-up studies.
+2 Evolution of polarization states in strong magnetic field – dipole pulsars
 2.1 Evolution equations of Stokes vector
-Classically, polarization properties of pulsar emission are described by the Stokes parameters $\left\{I,Q,U,V\right\}$, where $I$ is a measure of the total intensity, $Q$ and $U$ jointly describe the linear polarization, and $V$ describes the circular polarization of the curvature emission in pulsars (for more details, see Appendix A).
-However, in the presence of strong magnetic fields in the background of the emission, the polarization evolves along the photon propagation path from the emission point towards an observer.
-The evolution of the polarization can be investigated systematically using the formalism initiated by Kubo & Nagata (1981, 1983, 1985), and further developed by Heyl & Shaviv (2000), Heyl et al.
-(2003), Heyl & Caiazzo (2018), and Novak et al.
-(2018), namely, a system of evolution equations of the Stokes vector, described as
+Classically, polarization properties of pulsar emission are described by the Stokes parameters $\left\{I,Q,U,V\right\}$, where $I$ is a measure of the total intensity, $Q$ and $U$ jointly describe the linear polarization, and $V$ describes the circular polarization of pulsar emission (for more details, see Appendix A).
+However, in the presence of a strong magnetic field in the background of the emission, the polarization evolves along the photon propagation path from the emission point towards an observer.
+The evolution of the polarization can be investigated systematically using the formalism initiated by Kubo1981 ; Kubo1983 ; Kubo1985 , and further developed by Heyl2000 ; Heyl2003highenergy ; Heyl2018Strongly ; Novak2018 , namely, a system of evolution equations of the Stokes vector, described as
 $\frac{\mathrm{d}\mathbf{S}}{\mathrm{d}s}=k\mathbf{\hat{\Omega}}\times\mathbf{S},$ (1)
 where $k\equiv\omega/c$ denotes the wave number for the electromagnetic radiation and $s$ is an affine parameter to measure the length of the photon trajectory, and $\mathbf{S}$ is the normalized Stokes vector, defined out of the Stokes parameters as $\mathbf{S}=\left(S_{1},S_{2},S_{3}\right)\allowbreak\equiv\left(Q/I,U/I,V/I\right)$, [^1] The classical Stokes vector can be expressed via pulse profiles of pulsar curvature emission, as illustrated in Appendix A.
-and $\mathbf{\hat{\Omega}}$ is the birefringent vector, defined as [^2] Note that our $k\mathbf{\hat{\Omega}}$ is equivalent to the birefringent vector as defined in the references above.
+and $\mathbf{\hat{\Omega}}$ is the dimensionless birefringent vector, defined as [^2] Note that our $k\mathbf{\hat{\Omega}}$ is equivalent to the birefringent vector as defined in the references above.
 $\mathbf{\hat{\Omega}}\equiv\frac{\alpha_{\mathrm{e}}}{30\pi}\left(B/B_{\mathrm{c}}\right)^{2}\sin^{2}\vartheta\left(\mathcal{E}_{\mathrm{I}}^{2}-\mathcal{E}_{\mathrm{II}}^{2},2\mathcal{E}_{\mathrm{I}}\mathcal{E}_{\mathrm{II}},0\right),$ (2)
@@
 $\mathcal{E}_{i}\equiv-\mathbf{\hat{B}\cdot}\left(\mathbf{\hat{n}}_{\left[0\right]}\times\boldsymbol{\varepsilon}_{i\left[0\right]}\right),~{}~{}i=\mathrm{I},\mathrm{II},$ (4)
-with $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ being the two classical mode polarization vectors, orthogonal to each other and to $\mathbf{\hat{n}}_{\left[0\right]}$; the specific forms of $\mathbf{\hat{n}}_{\left[0\right]}$, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ are later given by equations (8), (12) and (13), respectively, for the magnetic field of an oblique dipole rotator as described by equation (5) and Fig.
+with $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ being the two classical mode polarization vectors, orthogonal to each other and to $\mathbf{\hat{n}}_{\left[0\right]}$; the specific forms of $\mathbf{\hat{n}}_{\left[0\right]}$, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$ are later given by Eqs.
+(8), (12) and (13), respectively, for the magnetic field of an oblique dipole rotator as described by Eq.
+(5) and Fig.
 1.
@@
 The scale of the unity in this graph is equivalent to the neutron star radius $\sim 10^{6}\,\mathrm{cm}$.
-The red dotted line represents the trajectory curve of the light ray traced by $\mathbf{\hat{n}}_{\left[0\right]}$ as projected onto the $xz$-plane.
-(Credit: Kim & Trippe (2021), reproduced with modifications.)
+The red dashed line represents the trajectory curve of the light ray traced by the propagation vector $\mathbf{\hat{n}}_{\left[0\right]}$ as projected onto the $xz$-plane.
+(Credit: Kim2021Generala , reproduced with modifications.)
 In our pulsar emission model, we consider curvature radiation produced along the magnetic field lines of an oblique dipole rotator as illustrated in Fig. 1:
-$\mathbf{B}\left(r,\theta,\phi\right)=\frac{2\mu\left(\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi\right)}{r^{3}}\mathbf{e}_{\hat{r}}+\frac{\mu\left(\cos\alpha\sin\theta-\sin\alpha\cos\theta\cos\phi\right)}{r^{3}}\mathbf{e}_{\hat{\theta}}+\frac{\mu\sin\alpha\sin\phi\,}{r^{3}}\mathbf{e}_{\hat{\phi}},$ (5)
+$\mathbf{B}\left(r,\theta,\phi\right)$ $=\frac{2\mu\left(\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi\right)}{r^{3}}\mathbf{e}_{\hat{r}}+\frac{\mu\left(\cos\alpha\sin\theta-\sin\alpha\cos\theta\cos\phi\right)}{r^{3}}\mathbf{e}_{\hat{\theta}}$
+$\hskip 12.0pt+\frac{\mu\sin\alpha\sin\phi\,}{r^{3}}\mathbf{e}_{\hat{\phi}},$ (5)
 where $\mu$ is the magnetic dipole moment and $\alpha$ denotes the inclination angle between the rotation axis and the magnetic axis.
@@
 The photon beam from curvature radiation is tangent to the field line at the emission point $\left(x_{\mathrm{o}},y_{\mathrm{o}},z_{\mathrm{o}}\right)=\left(r_{\mathrm{o}}\sin\theta_{\mathrm{o}},0,r_{\mathrm{o}}\cos\theta_{\mathrm{o}}\right)$.
-However, at the same time, our pulsar magnetosphere rotates, and therefore the field lines get twisted due to the magneto-centrifugal acceleration on the plasma particles moving along the field lines (Blandford & Payne (1982)).
-Then, taking into consideration this magneto-hydrodynamic (MHD) effect, the direction of the classical photon propagation, which must line up with the particle velocity in order for an observer to receive the radiation, can be described as (Gangadhara (2005))
-$\mathbf{\hat{n}}_{\left[0\right]}=\beta\mathbf{\hat{B}}+\mathbf{\Omega}_{\mathrm{p}}\times\mathbf{r,}$ (6)
+However, at the same time, our pulsar magnetosphere rotates, and therefore the field lines get twisted due to the magneto-centrifugal acceleration on the plasma particles moving along the field lines Blandford1982 .
+Then, taking into consideration this magneto-hydrodynamic (MHD) effect, the direction of the classical photon propagation, which must line up with the particle velocity in order for an observer to receive the radiation, can be described as Gangadhara2005
+$\mathbf{\hat{n}}_{\left[0\right]}=\beta\mathbf{\hat{B}}+\frac{\mathbf{\Omega}_{\mathrm{p}}\times\mathbf{r}}{c},$ (6)
 where on the right-hand side
@@
 In our analysis, the photon propagation is described with the consideration of the MHD effect above, assuming $\phi$ to be very small; e.g., $\phi\lesssim 10^{-1}$ is considered for a millisecond pulsar with $\Omega_{\mathrm{p}}\sim 10^{2}\,\mathrm{Hz}$, during the time of rotation $t\lesssim 10^{-3}\,\mathrm{s}$, such that $s\lesssim 10^{7}\,\mathrm{cm}$, which corresponds to the propagation distance within about $10$ times the neutron star radius.
-For equation (6) we take only the leading order expansions of $\mathbf{\hat{B}}\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ and $\beta\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ in $\phi$ from equations (5) and (7), respectively, and can express the classical propagation vector $\mathbf{\hat{n}}_{\left[0\right]}$ in Cartesian coordinates as
+For Eq.
+(6) we take only the leading order expansions of $\mathbf{\hat{B}}\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ and $\beta\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ in $\phi$ from Eqs.
+(5) and (7), respectively, and can express the classical propagation vector $\mathbf{\hat{n}}_{\left[0\right]}$ in Cartesian coordinates as
 $\mathbf{\hat{n}}_{\left[0\right]}=\hat{n}_{x\left[0\right]}\mathbf{e}_{x}+\hat{n}_{y\left[0\right]}\mathbf{e}_{y}+\hat{n}_{z\left[0\right]}\mathbf{e}_{z}\$ (8)
@@
 $\hat{n}_{y\left[0\right]}\approx\frac{\Omega_{\mathrm{p}}}{c}\left[\frac{\sin\alpha\,s}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}}+r_{\mathrm{o}}\sin\theta_{\mathrm{o}}\right]+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (11)
-where we have considered $\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\lesssim\phi$, e.g., for a millisecond pulsar with $\Omega_{\mathrm{p}}\sim 10^{2}\,\mathrm{Hz}$ and $r_{\mathrm{o}}\sim 10^{6}\,\mathrm{cm}$, such that $\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\lesssim\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\lesssim\phi^{2}$, all to be ignored in our analysis, and have substituted $\phi=\Omega_{\mathrm{p}}s/c$ in equation (11), the leading order rotational effect to be considered in our analysis.
-The orthogonal pair of classical mode polarization vectors, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$, both being also orthogonal to $\mathbf{\hat{n}}_{\left[0\right]}$ as given by equation (8) above, are determined as
+where we have considered $\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\lesssim\phi$, e.g., for a millisecond pulsar with $\Omega_{\mathrm{p}}\sim 10^{2}\,\mathrm{Hz}$ and $r_{\mathrm{o}}\sim 10^{6}\,\mathrm{cm}$, such that $\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\lesssim\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\lesssim\phi^{2}$, all to be ignored in our analysis, and have substituted $\phi=\Omega_{\mathrm{p}}s/c$ in Eq.
+(11), the leading order rotational effect to be considered in our analysis.
+The orthogonal pair of classical mode polarization vectors, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}$, both being also orthogonal to $\mathbf{\hat{n}}_{\left[0\right]}$ as given by Eq.
+(8) above, are determined as
 $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}$ $=\hat{n}_{z\left[0\right]}\mathbf{e}_{x}+\hat{n}_{y\left[0\right]}\mathbf{e}_{y}-\hat{n}_{x\left[0\right]}\mathbf{e}_{z},$ (12)
@@
 [^5] It can be checked out that $\mathbf{\hat{n}}_{\left[0\right]}\cdot\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}\approx 0+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$, $\mathbf{\hat{n}}_{\left[0\right]}\cdot\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}=0$ and $\boldsymbol{\varepsilon}_{\,\mathrm{I}\left[0\right]}\cdot\boldsymbol{\ \varepsilon}_{\mathrm{II}\left[0\right]}=0$ while $\mathbf{\hat{n}}_{\left[0\right]}^{2}\approx 1+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$ and $\boldsymbol{\ \varepsilon}_{\mathrm{I},\mathrm{II}\left[0\right]}^{2}\approx 1+\mathcal{O}\left(\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2}\right)$.
-Using these for equation (4), we obtain
+Using these for Eq.
+(4), we obtain
 $\mathcal{E}_{\mathrm{I}}$ $=-\mathbf{\hat{B}\cdot}\left(\mathbf{\hat{n}}_{\left[0\right]}\times\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]}\right)$
-$\approx\frac{4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)+2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}$
-$\hskip 10.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (14)
+$\approx\frac{4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)+2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\hat{n}_{y\left[0\right]}$
+$\hskip 12.0pt-\frac{\Omega_{\mathrm{p}}\sin\alpha\,s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (14)
 $\mathcal{E}_{\mathrm{II}}$ $=-\mathbf{\hat{B}\cdot}\left(\mathbf{\hat{n}}_{\left[0\right]}\times\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]}\right)$
-$\approx-\frac{2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right).$ (15)
-By means of equations (3), (5) and (8) one can express
-$\cos\vartheta$ $\approx\frac{4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (16)
-$\sin\vartheta$ $\approx\frac{2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right).$ (17)
-Now, using the relations between equations (14), (15) and (16), (17), the birefringent vector can finally be specified from equation (2):
+$\approx-\frac{2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}$
+$\hskip 12.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right).$ (15)
+By means of Eqs. (3), (5) and (8) one can express
+$\cos\vartheta$ $\approx\frac{4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}$
+$\hskip 12.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (16)
+$\sin\vartheta$ $\approx\frac{2\sin\left(\theta-\theta_{\mathrm{o}}\right)}{\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)+1\right)^{1/2}\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}$
+$\hskip 12.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right).$ (17)
+Now, using the relations between Eqs. (14)-(17), the birefringent vector can finally be specified from Eq. (2):
 $\hat{\Omega}_{1}$ $\approx-\eta B^{2}\sin^{4}\vartheta+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (18)
-$\hat{\Omega}_{2}$ $\approx 2\eta B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (19)
-where $\eta\equiv\alpha_{\mathrm{e}}/\left(30\pi B_{\mathrm{c}}^{2}\right)$ and
+$\hat{\Omega}_{2}$ $\approx-2\eta B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha\,s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]$
+$\hskip 12.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (19)
+where $\eta\equiv\alpha_{\mathrm{e}}/\left(30\pi B_{\mathrm{c}}^{2}\right)$666 $\eta=\eta_{2}-\eta_{1}$, where $\eta_{1}$ and $\eta_{2}$ are parameters defined via $\eta_{1}/4=\eta_{2}/7=\alpha_{\mathrm{e}}/\left(90\pi B_{\mathrm{c}}^{2}\right)\sim 10^{-31}\,\mathrm{g}^{-1}\,\mathrm{cm}\,\mathrm{s}^{2}$, from the post-Maxwellian Lagrangian $\mathcal{L}_{\mathrm{PM}}=-\left(\mathbf{B}^{2}-\mathbf{E}^{2}\right)/2+\eta_{1}\left(\mathbf{B}^{2}-\mathbf{E}^{2}\right)^{2}/4+\eta_{2}\left(\mathbf{E}\cdot\mathbf{B}\right)^{2}$ Euler1935Ueber .
+and
 $B=\frac{B_{\mathrm{max}}r_{*}^{3}\ \left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}{2\left(x^{2}+z^{2}\right)^{3/2}},$ (20)
-with $B_{\mathrm{max}}$ being the maximum magnetic field strength at the polar cap [^6] From equation (5) $B_{\mathrm{max}}=\left|B\left(r=r_{*},\theta=\alpha\right)\right|$.
-and $r_{*}$ being the neutron star radius ($\approx 10^{6}\,\mathrm{cm}$), and $\hat{n}_{y\left[0\right]}$, $\cos\vartheta$ and $\sin\vartheta$ are given by equations (11), (16) and (17), respectively.
-To facilitate solving the evolution equations (1) in the next Subsection, we substitute the following identities,
-$\cos\left(\theta-\alpha\right)=\frac{\sin\alpha x+\cos\alpha z}{\left(x^{2}+z^{2}\right)^{1/2}},~{}~{}\sin\left(\theta-\alpha\right)=\frac{\cos\alpha x-\sin\alpha z}{\left(x^{2}+z^{2}\right)^{1/2}},~{}~{}\sin\left(\theta-\theta_{\mathrm{o}}\right)=\frac{\cos\theta_{\mathrm{o}}x-\sin\theta_{\mathrm{o}}z}{\left(x^{2}+z^{2}\right)^{1/2}},$ (21)
+with $B_{\mathrm{max}}$ being the maximum magnetic field intensity at the polar cap [^7] From Eq.
+(5) $B_{\mathrm{max}}=\left|B\left(r=r_{*},\theta=\alpha\right)\right|$.
+and $r_{*}$ being the neutron star radius ($\approx 10^{6}\,\mathrm{cm}$), and $\hat{n}_{y\left[0\right]}$, $\cos\vartheta$ and $\sin\vartheta$ are given by Eqs.
+(11), (16) and (17), respectively.
+To facilitate solving the evolution equation (1) in the next subsection, we substitute the following identities,
+$\cos\left(\theta-\alpha\right)=\frac{\sin\alpha\,x+\cos\alpha\,z}{\left(x^{2}+z^{2}\right)^{1/2}},~{}\sin\left(\theta-\alpha\right)=\frac{\cos\alpha\,x-\sin\alpha\,z}{\left(x^{2}+z^{2}\right)^{1/2}},$
+$\sin\left(\theta-\theta_{\mathrm{o}}\right)=\frac{\cos\theta_{\mathrm{o}}\,x-\sin\theta_{\mathrm{o}}\,z}{\left(x^{2}+z^{2}\right)^{1/2}},$ (21)
 together with
-$x=\hat{n}_{x\left[0\right]}s+r_{\mathrm{o}}\sin\theta_{\mathrm{o}},~{}~{}z=\hat{n}_{z\left[0\right]}s+r_{\mathrm{o}}\cos\theta_{\mathrm{o}}$ (22)
-into (16) and (17). Then our solutions for the Stokes vector $\mathbf{S}$ will be parametrised by $s$.
-2.2 Solving evolution equations
-From equation (1) we write down a system of first-order ordinary differential equations to solve:
+$x=\hat{n}_{x\left[0\right]}s+r_{\mathrm{o}}\sin\theta_{\mathrm{o}},~{}z=\hat{n}_{z\left[0\right]}s+r_{\mathrm{o}}\cos\theta_{\mathrm{o}}$ (22)
+into (16) and (17). Then our solutions for the Stokes vector $\mathbf{S}$ will be parameterized solely by $s$.
+2.2 Solving the evolution equations
+From Eq. (1) we write down a system of first-order ordinary differential equations to solve:
 $\dot{S}_{1}\left(s\right)$ $=k\hat{\Omega}_{2}\left(s\right)S_{3}\left(s\right),$ (23)
@@
 By solving these equations numerically, we find out how the photon polarization evolves through the strong magnetic field in the background of our pulsar emission.
-However, in case $\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\ll 1$, one can obtain a solution to equation (1) via perturbation:
+However, in case $\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\ll 1$, one can obtain a solution to Eq. (1) via perturbation:
 $\mathbf{S}=\mathbf{S}_{\left[0\right]}+\delta\mathbf{S}_{\left[1\right]}=\mathbf{S}_{\left[0\right]}+k\int\mathbf{\hat{\Omega}}\times\mathbf{S}_{\left[0\right]}\,\mathrm{d}s,$ (26)
@@
 Here the correction can be treated as the leading order perturbation with $\alpha_{\mathrm{e}}/\left(30\pi\right)\left(B/B_{\mathrm{c}}\right)^{2}\sim 10^{-5}\left(B/B_{\mathrm{c}}\right)^{2}$ being a perturbation parameter.
-Upon inspection of equations (18) and (19) for equation (26), we can further write down our solution in terms of its components:
-$S_{1}$ $\approx S_{1\left[0\right]}+2k\eta S_{3\left[0\right]}\int B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]\mathrm{d}s,$ (27)
+Upon inspection of Eqs.
+(18) and (19) for Eq.
+(26), we can further write down our solution in terms of its components:
+$S_{1}$ $\approx S_{1\left[0\right]}-2k\eta S_{3\left[0\right]}\int B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha\,s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]\mathrm{d}s,$ (27)
 $S_{2}$ $\approx S_{2\left[0\right]}+k\eta S_{3\left[0\right]}\int B^{2}\sin^{4}\vartheta\mathrm{d}s,$ (28)
-$S_{3}$ $\approx S_{3\left[0\right]}-k\eta\left\{S_{2\left[0\right]}\int B^{2}\sin^{4}\vartheta\mathrm{d}s+2S_{1\left[0\right]}\int B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]\mathrm{d}s\right\}.$ (29)
+$S_{3}$ $\approx S_{3\left[0\right]}+k\eta\left\{2S_{1\left[0\right]}\int B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha\,s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]\mathrm{d}s\right.$
+$\!\left.\left.\hskip 12.0pt-S_{2\left[0\right]}\int B^{2}\sin^{4}\vartheta\mathrm{d}s\right._{\,{}_{\,{}_{\,}}}^{\,{}^{\,{}^{\,}}}\!\!\!\!\!\right\}.$ (29)
 2.2.1 Examples
-We consider X-ray emissions from three neutron stars, (i) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (ii) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (iii) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), and assume for all the emissions, $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$).
-These stars belong to ‘rotation-powered pulsars’ (RPPs) (Pavlov et al.
-(2013)).
+We consider X-ray emissions from three neutron stars: (i) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (ii) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (iii) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$).
+For all three, we assume $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$).
+[^8] For the emission location $\left(r_{\mathrm{o}},\theta_{\mathrm{o}}\right)$ and the inclination angle $\alpha$ are given the same values for the three stars; the values are not based on actual observations.
+This is intended for comparing the QED effects from the three different sources under the same conditions.
+These stars belong to ‘rotation-powered pulsars’ (RPPs) Pavlov2013 .
 [^9] RPPs refer to neutron stars whose radiation is powered by loss of their rotation energy, via creation and acceleration of $e^{+}e^{-}$ pairs in the strong magnetic field, $B_{\mathrm{max}}\sim 10^{11}-10^{13}\,\mathrm{G}$.
-The number of detected RPPs are known to be about $\sim 2000$ in radio, $\sim 10$ in optical (including NIR and UV), $\sim 100$ in X-ray and $\sim 100$ in gamma-ray emissions (Pavlov et al.
-(2013)).
+The number of detected RPPs are known to be about $\sim 4000$ in radio, $\sim 10$ in optical (including NIR and UV), $\sim 100$ in X-ray and $\sim 300$ in gamma-ray emissions Pavlov2013 ; ANTF2024 ; Smith_2023 .
 In Fig.
-2 the three RPPs chosen from the X-ray group are encircled, (i) the one in orange, (ii) another in cyan, (iii) the third in green.
+2 the three RPPs chosen from the X-ray group are encircled: (i) the one in orange, (ii) another in cyan, (iii) the third in green.
 Given the X-ray emissions from these, we solve the evolution equations (1) for the following cases.
@@
 RPPs detected in X-rays and gamma-rays are represented by red dots and blue stars, respectively.
-Three RPPs chosen from the X-ray group, (i) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (ii) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (iii) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), are encircled in orange, cyan and green colours, respectively.
-(Credit: Pavlov et al.
-(2013), reproduced with modifications.)
+Three RPPs chosen from the X-ray group, (i) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (ii) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (iii) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), are encircled in orange, cyan and green colors, respectively.
+(Credit: Pavlov2013 , reproduced with modifications.)
 Example (i)
-With $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), we obtain numerical solutions through equations (23)-(25) (in solid lines) or perturbative solutions by means of equations (27)-(29) (in dashed lines), as shown in Figs.
+With $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), we obtain numerical solutions of Eqs.
+(23)-(25) (in solid lines) or perturbative solutions by means of Eqs.
+(27)-(29) (in dashed lines), as shown in Figs.
 3a and 3b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively.
-The perturbative solutions agree well with numerical ones as $\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\sim\left|k\hat{\Omega}_{1,2\left(\mathrm{min}\right)}s_{\dagger 1,2}\right|\sim 10^{-2}\ll 1$, where $\hat{\Omega}_{1,2\left(\mathrm{min}\right)}\equiv\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)$, evaluated from the birefringent functions given by (18), (19), with $s_{\dagger 1,2}$ being the extremum, i.e., $\mathrm{d}\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)/\mathrm{d}s=0$.
+The perturbative solutions agree well with numerical ones as $\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\sim\left|k\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)s_{\dagger 1,2}\right|\sim 10^{-2}\ll 1$, where $s_{\dagger 1,2}$ is the extremum, i.e., $\mathrm{d}\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)/\mathrm{d}s=0$.
 On the Poincaré sphere, our solutions are represented by the magenta and light blue loci in Fig.
@@
 3a and 3b, respectively.
-The loci imply a fraction of an oscillation for the polarization evolution, as is confirmed later by the approximate analytical solutions in Section 2.2.2.
+The loci imply a fraction of an oscillation for the polarization evolution, as is confirmed later by the approximate analytical solutions in Sect.
+2.2.2.
 (a) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$
 (b) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$
-Figure 3: Corresponding to Example (i): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$).
+Figure 3: For Example (i): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$); the subscript $\left(\mathrm{p}\right)$ stands for ‘perturbative’.
 Example (ii)
-With $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), we obtain numerical solutions through equations (23)-(25), as shown in Figs.
+With $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), we obtain numerical solutions of Eqs.
+(23)-(25), as shown in Figs.
 4a and 4b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively.
@@
 4a and 4b, respectively.
-The loci imply about half an oscillation for the polarization evolution, as is confirmed later by the approximate analytical solutions in Section 2.2.2.
+The loci imply about half an oscillation for the polarization evolution, as is confirmed later by the approximate analytical solutions in Sect.
+2.2.2.
 (a) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$
 (b) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$
-Figure 4: Corresponding to Example (ii): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$).
+Figure 4: For Example (ii): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$).
 Example (iii)
-With $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), we obtain numerical solutions through equations (23)-(25), as shown in Figs.
+With $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), we obtain numerical solutions of Eqs.
+(23)-(25), as shown in Figs.
 5a and 5b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively.
@@
 5a and 5b, respectively.
-The loci imply multiple oscillations for the polarization evolution, as is confirmed later by the approximate analytical solutions in Section 2.2.2.
+The loci imply multiple oscillations for the polarization evolution, as is confirmed later by the approximate analytical solutions in Sect.
+2.2.2.
 (a) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$
 (b) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$
-Figure 5: Corresponding to Example (iii): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$).
-(a) Corresponding to Example (i):
-a fraction of an oscillation
-(b) Corresponding to Example (ii):
-about half an oscillation
-(c) Corresponding to Example (iii):
-multiple oscillations
-Figure 6: Representations of the Stokes vectors on the Poincaré sphere. The loci imply patterns of the polarization evolution in terms of oscillation.
+Figure 5: For Example (iii): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$).
+(a) Example (i)
+(b) Example (ii)
+(c) Example (iii)
+Figure 6: Representations of the Stokes vectors from Examples (i)-(iii) on the Poincaré sphere. The loci imply patterns of the polarization evolution in terms of oscillation.
+With regard to the adiabatic evolution condition as mentioned in Refs.
+Heyl2000 ; Heyl2003highenergy ; Heyl2018Strongly , we carefully examine our results presented in Figs.
+3-5 to see what interpretations the condition leads to.
+Solving the condition $\left|k\mathbf{\hat{\Omega}}\left(\mathrm{d}\ln|k\mathbf{\hat{\Omega}}|/\mathrm{d}s\right)^{-1}\right|\gtrsim 0.05$ for $s$ yields $s_{\mathrm{PL}1}\lesssim s\lesssim s_{\mathrm{PL}2}$, [^10] Note that our $k\mathbf{\hat{\Omega}}$ is equivalent to the birefringent vector as defined in Refs.
+Heyl2000 ; Heyl2003highenergy ; Heyl2018Strongly and that we set the condition value to $0.05$ rather than $0.5$ as in the references.
+where $s_{\mathrm{PL}1[2]}$ refers to the lower [upper] bound for the ‘polarization limiting’ distance as measured from the emission point.
+Using this, one can check out the following:
+(1) for $6.2\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 1.8\times 10^{6}\,\mathrm{cm}$ in Fig. 3,
+(2) for $3.1\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 6.2\times 10^{6}\,\mathrm{cm}$ in Fig. 4,
+(3) for $1.1\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 1.9\times 10^{7}\,\mathrm{cm}$ in Fig. 5,
+our Stokes vector evolves evidently; otherwise, it freezes.
 2.2.2 Approximate analytical solutions
@@
 7 are plotted the birefringent functions for the three cases: (a) $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (b) $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (c) $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$) with solid lines (see Figs.
-7a, 7b and 7c, respectively), where they have been evaluated with the same initial condition as assumed in Section 2.2.1.
+7a, 7b and 7c, respectively), where they have been evaluated with the same initial condition as assumed in Sect.
+2.2.1.
 In correspondence with the actual birefringent functions above, the following analytical models are also plotted with dashed lines in Fig. 7:
 $\hat{\Omega}_{1,2}\left(s\right)\approx-a_{1,2}s^{\frac{p+1}{p}}e^{-bs}~{}~{}\text{for $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$},$ (30)
-where $a_{1,2}>0$, $b>0$ and $p>0$ are free parameters; with suitable values chosen for these, our model functions can give rise to solutions of equations (23)-(25) that match fairly well the numerical results obtained in Section 2.2.1.
+where $a_{1}>0$, $a_{2}<0$, $b>0$ and $p>0$ are free parameters; with suitable values chosen for these, our model functions can give rise to solutions of Eqs.
+(23)-(25) that match fairly well the numerical results obtained in Sect.
+2.2.1.
 Here we express
-$a_{1,2}=-\hat{\Omega}_{1,2\left(\mathrm{min}\right)}\left[\frac{e}{\left(1-\chi\right)s_{\dagger 1}+\chi s_{\dagger 2}}\right]^{\frac{p+1}{p}},~{}~{}b=\frac{p+1}{p\left[\left(1-\chi\right)s_{\dagger 1}+\chi s_{\dagger 2}\right]},$ (31)
-where $\hat{\Omega}_{1,2\left(\mathrm{min}\right)}=\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)$, evaluated from (18), (19), with $s_{\dagger 1,2}$ denoting the extremum.
-We have set $p=100$ (a sufficiently large number) for all three cases, and $\chi=0.999$ for (a) and (b), and $\chi=0.056$ for (c) in Fig.
-7.888In particular, the values for $\chi$ have been chosen such that our solutions converge to the asymptotic limits that match well the numerical results given by Figs.
-3b, 4b and 5b in Section 2.2.1, as $s$ tends to $\infty$.
+$a_{1,2}=-\hat{\Omega}_{1\left(\mathrm{min}\right),2\left(\mathrm{max}\right)}\left[\frac{e}{\left(1-q\right)s_{\dagger 1}+qs_{\dagger 2}}\right]^{\frac{p+1}{p}},~{}~{}b=\frac{p+1}{p\left[\left(1-q\right)s_{\dagger 1}+qs_{\dagger 2}\right]},$ (31)
+where $\hat{\Omega}_{1\left(\mathrm{min}\right),2\left(\mathrm{max}\right)}=\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)$, evaluated from (18), (19), with $s_{\dagger 1,2}$ denoting the extremum.
+We have set $p=100$ (a sufficiently large number) for all three cases, and $q=0.99999$ for (a) and (b), and $q=0.0532$ for (c) in Fig.
+7.111111In particular, the values for $q$ have been chosen such that our solutions converge to the asymptotic limits that match well the numerical results given by Figs.
+3b, 4b and 5b in Sect.
+2.2.1, as $s$ tends to $\infty$.
 (a) For $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$)
@@
 (c) For $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$)
-Figure 7: Plots of the birefringent functions $\hat{\Omega}_{1}\left(s\right)$ and $\hat{\Omega}_{2}\left(s\right)$ and their approximate analytic models.
-Plugging equation (30) into equations (23)-(25), we obtain analytical solutions as follows (for a complete derivation, see Appendix B):
+Figure 7: Plots of the birefringent functions $\hat{\Omega}_{1}\left(s\right)$ and $\hat{\Omega}_{2}\left(s\right)$ and their approximate analytic models $\hat{\Omega}_{1\left(\mathrm{m}\right)}\left(s\right)$ and $\hat{\Omega}_{2\left(\mathrm{m}\right)}\left(s\right)$.
+Plugging Eq. (30) into Eqs. (23)-(25), we obtain analytical solutions as follows (for a complete derivation, see Appendix B):
 $S_{1}\left(s\right)$ $\approx\frac{a_{2}S_{\mathrm{o}}}{\sqrt{a_{1}^{2}+a_{2}^{2}}}\cos\left(\Psi\left(s;p\right)+\delta\right)+\frac{a_{1}C}{\sqrt{a_{1}^{2}+a_{2}^{2}}},$ (32)
@@
 $\Psi\left(s;p\right)\equiv k\sqrt{a_{1}^{2}+a_{2}^{2}}b^{-\frac{4p+1}{2p}}s^{\frac{1}{2p}}e^{-\frac{1}{2}bs}\left[M_{\frac{1}{2p},\frac{p+1}{2p}}\left(bs\right)-\left(bs\right)^{\frac{2p+1}{2p}}e^{-\frac{1}{2}bs}\right],$ (35)
-and $M_{\kappa,\mu}\left(z\right)$ denotes a Whittaker function of the first kind, [^9] The expression inside the square brackets in equation (35) has been reduced from its original form as given by equation (60) in Appendix B, using the identity $M_{\left(2p+1\right)/\left(2p\right),\left(p+1\right)/\left(2p\right)}\left(bs\right)=\left(bs\right)^{\left(2p+1\right)/\left(2p\right)}e^{-bs/2}M\left(0,\left(2p+1\right)/p,bs\right)$, with the Kummer function $M\left(0,\left(2p+1\right)/p,bs\right)=1$ as a special case (NIST (2024)).
+and $M_{\kappa,\mu}\left(z\right)$ denotes a Whittaker function of the first kind, [^12] The expression inside the square brackets in Eq.
+(35) has been reduced from its original form as given by Eq.
+(83) in Appendix B, using the identity $M_{\left(2p+1\right)/\left(2p\right),\left(p+1\right)/\left(2p\right)}\left(bs\right)=\left(bs\right)^{\left(2p+1\right)/\left(2p\right)}e^{-bs/2}M\left(0,\left(2p+1\right)/p,bs\right)$, with the Kummer function $M\left(0,\left(2p+1\right)/p,bs\right)=1$ as a special case NIST2024 .
 and $a_{1,2}$ and $b$ are given by (31).
-Here we determine $S_{\mathrm{o}}$, $C$ and $\delta$ by matching the initial value of the Stokes vector $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)$ with equations (32)-(34) evaluated at $s=0$.
+Here we determine $S_{\mathrm{o}}$, $C$ and $\delta$ by matching the initial value of the Stokes vector $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)$ with Eqs.
+(32)-(34) evaluated at $s=0$.
 In Fig.
-8 we plot the above solutions for the following cases, assuming $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ for $\hat{\Omega}_{1,2\left(\mathrm{min}\right)}$ (evaluated through (18), (19)) and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$) for the X-ray emissions, with the initial Stokes vector $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$:
+8 we plot the above solutions (32)-(34) for the following cases, assuming $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ for $\hat{\Omega}_{1\left(\mathrm{min}\right),2\left(\mathrm{max}\right)}=\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)$ (evaluated via (18) and (19)) and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$) for the X-ray emissions, with the initial Stokes vector $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$:
 Case (a) (see Fig. 8a);
 for $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), with the parameters
-$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-2.5511\times 10^{-16}$, $\hat{\Omega}_{2\left(\mathrm{min}\right)}\approx-2.2054\times 10^{-16}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $\chi=0.999$, $S_{\mathrm{o}}\approx 0.7960$, $C\approx 0.6052$ and $\delta\approx 0.8537$,
+$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-2.5511\times 10^{-16}$, $\hat{\Omega}_{2\left(\mathrm{max}\right)}\approx 2.2054\times 10^{-16}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $q=0.99999$ for $a_{1,2}$ and $b$ (evaluated via (31)), and $S_{\mathrm{o}}\approx 0.7961$, $C\approx 0.6052$, $\delta\approx 2.2879$,
 Case (b) (see Fig. 8b);
 for $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), with the parameters
-$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-8.0001\times 10^{-15}$, $\hat{\Omega}_{2\left(\mathrm{min}\right)}\approx-3.9240\times 10^{-16}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $\chi=0.999$, $S_{\mathrm{o}}\approx 0.6013$, $C\approx 0.7990$ and $\delta\approx 1.5056$,
+$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-8.0001\times 10^{-15}$, $\hat{\Omega}_{2\left(\mathrm{max}\right)}\approx 3.9240\times 10^{-16}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $q=0.99999$ for $a_{1,2}$ and $b$ (evaluated via (31)), and $S_{\mathrm{o}}\approx 0.6013$, $C\approx 0.7990$, $\delta\approx 1.6360$,
 Case (c) (see Fig. 8c);
 for $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), with the parameters
-$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-6.3776\times 10^{-13}$, $\hat{\Omega}_{2\left(\mathrm{min}\right)}\approx-2.7567\times 10^{-14}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $\chi=0.056$, $S_{\mathrm{o}}\approx 0.6010$, $C\approx 0.7993$ and $\delta\approx 1.5133$.
-These plots compare with Figs. 3b (for Example (i)), 4b (for Example (ii)) and 5b (for Example (iii)) in Section 2.2.1, respectively.
+$\hat{\Omega}_{1\left(\mathrm{min}\right)}\approx-6.3776\times 10^{-13}$, $\hat{\Omega}_{2\left(\mathrm{max}\right)}\approx 2.7567\times 10^{-14}$, $s_{\dagger 1}\approx 1.0202\times 10^{6}\,\mathrm{cm}$, $s_{\dagger 2}\approx 7.5467\times 10^{5}\,\mathrm{cm}$, $p=100$, $q=0.0532$ for $a_{1,2}$ and $b$ (evaluated via (31)), and $S_{\mathrm{o}}\approx 0.6010$, $C\approx 0.7993$, $\delta\approx 1.6283$.
+These plots compare with Figs. 3b (for Example (i)), 4b (for Example (ii)) and 5b (for Example (iii)) in Sect. 2.2.1, respectively.
 (a) For $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$)
@@
 (c) For $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$)
-Figure 8: Plots of the analytical solutions $S_{1}\left(s\right)$, $S_{2}\left(s\right)$ and $S_{3}\left(s\right)$, given the initial Stokes vector $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$.
+Figure 8: Plots of the analytical solutions $\mathbf{S}_{\left(\mathrm{a}\right)}\left(s\right)=\left(S_{1\left(\mathrm{a}\right)}\left(s\right),S_{2\left(\mathrm{a}\right)}\left(s\right),S_{3\left(\mathrm{a}\right)}\left(s\right)\right)$, given the initial Stokes vector $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$.
 The analytical solutions (32)-(34) provide a useful tool for understanding the different patterns of polarization evolution for the three cases above, as given by Figs.
@@
 For $0\leq s\lesssim 2.5\times 10^{6}\,\mathrm{cm}$,
-$\Psi\left(s;p=100\right)=k\sqrt{a_{1}^{2}+a_{2}^{2}}b^{-\frac{401}{200}}s^{\frac{1}{200}}e^{-\frac{1}{2}bs}\left[M_{\frac{1}{200},\frac{101}{200}}\left(bs\right)-\left(bs\right)^{\frac{201}{200}}e^{-\frac{1}{2}bs}\right]\approx 0.24k\sqrt{\hat{\Omega}_{1\left(\mathrm{min}\right)}^{2}+\hat{\Omega}_{2\left(\mathrm{min}\right)}^{2}}\pi s.$ (36)
-Using this, we can estimate how much the oscillations for the three cases have progressed, for example, during $0\leq s\lesssim 1.2\times 10^{6}\,\mathrm{cm}$:
-$\Psi\left(s\approx 1.2\times 10^{6}\,\mathrm{cm};p=100\right)$$\approx\left\{\begin{array}[]{ll}0.02\pi\hskip 9.0pt\text{(a fraction of an oscillation)}&\hskip 12.0pt\text{for Case (a)},\\ 0.5\pi\hskip 13.0pt\text{(about half an oscillation)}&\hskip 12.0pt\text{for Case (b)},\\ 39\pi\hskip 15.0pt\text{(multiple oscillations)}&\hskip 12.0pt\text{for Case (c)},\end{array}\right.$ (40)
-which can be checked by comparison with Figs. 8a, 8b and 8c, respectively.
-Recalling Example (i) from Section 2.2.1, one finds that the condition for perturbation can be equivalently expressed using (36):
+$\Psi\left(s;p=100\right)$ $=k\sqrt{a_{1}^{2}+a_{2}^{2}}b^{-\frac{401}{200}}s^{\frac{1}{200}}e^{-\frac{1}{2}bs}\left[M_{\frac{1}{200},\frac{101}{200}}\left(bs\right)-\left(bs\right)^{\frac{201}{200}}e^{-\frac{1}{2}bs}\right]$
+$\approx 0.24k\sqrt{\hat{\Omega}_{1\left(\mathrm{min}\right)}^{2}+\hat{\Omega}_{2\left(\mathrm{max}\right)}^{2}}\pi s.$ (36)
+Using this, we can estimate how much the oscillations for the three cases have progressed, for example, during $0\leq s\lesssim 2.5\times 10^{6}\,\mathrm{cm}$:
+$\Psi\left(s\approx 2.5\times 10^{6}\,\mathrm{cm};p=100\right)\approx\left\{\begin{array}[]{ll}0.04\pi\hskip 3.0pt\text{(a fraction of an oscillation)}&\hskip 3.0pt\text{for Case (a)},\\ \pi\hskip 8.0pt\text{(about half an oscillation)}&\hskip 3.0pt\text{for Case (b)},\\ 80\pi\hskip 11.0pt\text{(multiple oscillations)}&\hskip 3.0pt\text{for Case (c)},\end{array}\right.$ (40)
+each of which can be checked by comparison with Figs. 8a, 8b and 8c, respectively.
+Recalling Example (i) from Sect. 2.2.1, one finds that the condition for perturbation can be equivalently expressed with the help of (36):
+$\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\sim\left|k\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)s_{\dagger 1,2}\right|$ $\sim 0.24k\sqrt{\hat{\Omega}_{1\left(\mathrm{min}\right)}^{2}+\hat{\Omega}_{2\left(\mathrm{max}\right)}^{2}}\pi\times 10^{6}\,\mathrm{cm}$
 […]
 Extending this argument, similarly to (40) above we may state the following in reference to the patterns of polarization evolution:
-$\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\left\{\begin{array}[]{ll}\ll 1\hskip 9.0pt\text{(fractionally oscillatory - monotonic)}&\hskip 12.0pt\text{for Example (i)},\\ \sim\hskip 2.0pt1\hskip 9.0pt\text{(half oscillatory)}&\hskip 12.0pt\text{for Example (ii)},\\ \gg 1\hskip 9.0pt\text{(highly oscillatory)}&\hskip 12.0pt\text{for Example (iii)}.\end{array}\right.$ (45)
+$\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\left\{\begin{array}[]{ll}\ll 1\hskip 5.0pt\text{(fractionally oscillatory - monotonic)}&\hskip 3.0pt\text{for Example (i)},\\ \sim\hskip 2.0pt1\hskip 5.0pt\text{(half oscillatory)}&\hskip 3.0pt\text{for Example (ii)},\\ \gg 1\hskip 5.0pt\text{(highly oscillatory)}&\hskip 3.0pt\text{for Example (iii)}.\end{array}\right.$ (45)
+These features can also be checked by comparison with the loci on the Poincaré sphere, as given by Figs.
+6a, 6b and 6c, which imply the three different patterns of the polarization evolution in terms of oscillation (by means of the number of turns of the circular loci).
+Here the approximate analytical solutions serve our purpose well, in that they help us to understand the different patterns of polarization evolution, which depend largely on the major profiles of pulsar emission, such as the emission frequency, the magnetic field strength and the rotation frequency of the neutron stars, as implied from Eq.
+(36).
+On the other hand, it would be worthwhile to check how close the numerical and analytical solutions are to each other by evaluating the cross-correlations between them.
+For example, comparing Figs.
+5b and 8c, over the entire region of polarization evolution ($2\times 10^{2}\,\mathrm{cm}\lesssim s\lesssim 2\times 10^{7}\,\mathrm{cm}$), the cross-correlation coefficients between $S_{1}$ and $S_{1\left(\mathrm{a}\right)}$, between $S_{2}$ and $S_{2\left(\mathrm{a}\right)}$, and between $S_{3}$ and $S_{3\left(\mathrm{a}\right)}$ turn out to be approximately $0.9998$, $0.6757$, and $0.6521$, respectively.
+This shows that the two solutions are in decent agreement with each other; although the asymptotic values, $S_{1}\left(\infty\right)$ and $S_{1\left(\mathrm{a}\right)}\left(\infty\right)$, $S_{2}\left(\infty\right)$ and $S_{2\left(\mathrm{a}\right)}\left(\infty\right)$, and $S_{3}\left(\infty\right)$ and $S_{3\left(\mathrm{a}\right)}\left(\infty\right)$ are fairly closely matched to each other with less than $10\%$ differences, the correlations are decreased by the mismatched phases between the two solutions.
+3 Evolution of polarization states in strong magnetic field – quadrudipole pulsars
+3.1 Modified magnetic field geometry and evolution equations of Stokes vector
+The pulsar magnetic field structure may not be assumed to be purely dipolar as given by (5).
+This assumption is based on the behavior of the field in the far-field regime, where its high-order multipole ($\ell$) components decrease faster than low-order ones, like $r^{-\left(\ell+1\right)}$, which justifies the use of the dipole field as a good approximation Petri2016Theory .
+However, taking into account the possible contributions from the multipolar components, especially in the vicinity of the neutron star, we need to extend our pulsar model by superposing the dipole and higher-order multipole fields.
+A simple extension can be implemented by considering a rotating off-centered dipole, and a number of studies have been carried out regarding a variety of astrophysical consequences of such extension models in pulsar astronomy (see Petri2016Theory and references therein).
+[^13] The multipole field structure extended in this way inevitably has the higher-order fields aligned with the dipolar axis.
+Although the aligned multipole fields might not accurately represent actual field geometries in nature (as illustrated in Ref.
+Kazmierczak2019 ), the models would still be useful for estimating roughly the ‘beyond-dipole’ effects in pulsar emission in the near-field regime, as shown in Sect.
+3.2.1 later.
+As the simplest model, one can consider ‘quadrudipole’ fields, a superposition of dipole and quadrupole fields Gralla_2016 ; 10.1093/mnras/stz2524 ; Kalapotharakos_2021 .
+The magnetic field of an oblique quadrudipole rotator can be written as
+$\mathbf{B}_{\mathrm{qd}}\left(r,\theta,\phi\right)$
+$=\left[\frac{2\mu_{\mathrm{d}}\left(\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi\right)}{r^{3}}+\frac{\mu_{\mathrm{q}}\left(3\left(\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi\right)^{2}-1\right)}{r^{4}}\right]\mathbf{e}_{\hat{r}}$
+$\hskip 14.0pt+\left[\frac{\mu_{\mathrm{d}}\left(\cos\alpha\sin\theta-\sin\alpha\cos\theta\cos\phi\right)}{r^{3}}\right.$
+$\hskip 31.0pt\left.+\frac{2\mu_{\mathrm{q}}\left(\cos\alpha\sin\theta-\sin\alpha\cos\theta\cos\phi\right)\left(\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi\right)}{r^{4}}\right]\mathbf{e}_{\hat{\theta}}$
+$\hskip 14.0pt+\left[\frac{\mu_{\mathrm{d}}\sin\alpha\sin\phi}{r^{3}}+\frac{2\mu_{\mathrm{q}}\sin\alpha\sin\phi\left(\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi\right)}{r^{4}}\right]\mathbf{e}_{\hat{\phi}}.$ (46)
+Here the subscript ‘qd’ on the left-hand side stands for ‘quadrudipole’ (hereafter, this will be attached to notations for any quantities affected by the quadrudipole field), and $\mu_{\mathrm{d}}$ and $\mu_{\mathrm{q}}$ denote the magnetic dipole and quadrupole moments, respectively.
+The magnetic dipole moment can be expressed as $\mu_{\mathrm{d}}=\pi r_{\ast}^{2}I$, as produced by a static loop current $I$ of radius of the neutron star $r_{\ast}$, encircling its equator.
+Similarly, one can express the magnetic quadrupole moment as $\mu_{\mathrm{q}}=2\chi\pi r_{\ast}^{3}I=2\chi r_{\ast}\mu_{\mathrm{d}}$, as produced by two identical magnetic dipole loops carrying opposing equal currents $I$, each of radius $r_{\ast}$, separated by distance $\chi r_{\ast}$ (i.e., anti-Helmholtz coils), where $\chi>0$ is a free parameter to determine the ratio between the dipole and quadrupole moments.
+For example, the magnetic field lines of an oblique quadrudipole rotator, with $\chi=1$, that is, $\mu_{\mathrm{q}}=2r_{\ast}\mu_{\mathrm{d}}$, set for Eq.
+(46) are illustrated in Fig.
+9.
+Figure 9: A cross-sectional view of a pulsar magnetosphere with the quadrudipole (dipole $+$ quadrupole) magnetic field lines (green) around a neutron star, where the relation between the magnetic dipole and quadrupole moments is set by $\mu_{\mathrm{q}}=2r_{\ast}\mu_{\mathrm{d}}$ for the total field (46).
+The vertical dashed line (black) and the inclined solid line (red) represent the rotation axis and the magnetic axis, respectively.
+$\alpha$ between these axes denotes the inclination angle.
+The scale of the unity in this graph is equivalent to the neutron star radius $\sim 10^{6}\,\mathrm{cm}$.
+The red dashed line represents the trajectory curve of the light ray traced by the propagation vector $\mathbf{\hat{n}}_{\left[0\right]\mathrm{qd}}$ as projected onto the $xz$-plane.
+As the magnetic field geometry changes from (5) to (46), the classical propagation vector shall be modified from (6) to
+$\mathbf{\hat{n}}_{\left[0\right]\mathrm{qd}}=\beta_{\mathrm{qd}}\mathbf{\hat{B}}_{\mathrm{qd}}+\frac{\mathbf{\Omega}_{\mathrm{p}}\times\mathbf{r}}{c},$ (47)
+where $\mathbf{\hat{B}}_{\mathrm{qd}}\equiv\mathbf{B}_{\mathrm{qd}}/\left|\mathbf{B}_{\mathrm{qd}}\right|$ and
+$\beta_{\mathrm{qd}}\equiv\left[1-\left(\frac{\Omega_{\mathrm{p}}r}{c}\right)^{2}\sin^{2}\theta\left(1-\frac{\sin^{2}\alpha\sin^{2}\phi}{f\left(\theta^{\prime};\chi\right)}\right)\right]^{1/2}-\frac{\Omega_{\mathrm{p}}r}{c}\frac{\sin\alpha\sin\theta\sin\phi}{\left[f\left(\theta^{\prime};\chi\right)\right]^{1/2}},$ (48)
+with
+$f\left(\theta^{\prime};\chi\right)\equiv 3\cos^{2}\theta^{\prime}+1+16\chi\frac{r_{\ast}}{r}\cos^{3}\theta^{\prime}+4\chi^{2}\left(\frac{r_{\ast}}{r}\right)^{2}\left(5\cos^{4}\theta^{\prime}-2\cos^{2}\theta^{\prime}+1\right)$ (49)
+and $\cos\theta^{\prime}\equiv\cos\alpha\cos\theta+\sin\alpha\sin\theta\cos\phi$.
+Similarly to Sect.
+2.1, taking only the leading order expansions of $\mathbf{\hat{B}}_{\mathrm{qd}}\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ and $\beta_{\mathrm{qd}}\left(r_{\mathrm{o}},\theta_{\mathrm{o}},\phi\right)$ in $\phi$ from Eqs.
+(46) and (48), respectively, one can write down the classical propagation vector out of Eq.
+(47) in Cartesian coordinates as
+$\mathbf{\hat{n}}_{\left[0\right]\mathrm{qd}}=\hat{n}_{x\left[0\right]\mathrm{qd}}\mathbf{e}_{x}+\hat{n}_{y\left[0\right]\mathrm{qd}}\mathbf{e}_{y}+\hat{n}_{z\left[0\right]\mathrm{qd}}\mathbf{e}_{z}\$ (50)
+where
+$\hat{n}_{x\left[0\right]\mathrm{qd}}\approx$ $\left[f\left(\theta_{\mathrm{o}}-\alpha;\chi\right)\right]^{-1/2}\left\{2\cos\left(\theta_{\mathrm{o}}-\alpha\right)\sin\theta_{\mathrm{o}}+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\cos\theta_{\mathrm{o}}~{}_{~{}_{~{}_{~{}_{~}}}}^{~{}^{~{}^{~{}^{~}}}}\right.$
+$\left.+\hskip 1.0pt2\chi\frac{r_{\ast}}{r_{\mathrm{o}}}\left[\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)-1\right)\sin\theta_{\mathrm{o}}+2\cos\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta_{\mathrm{o}}-\alpha\right)\cos\theta_{\mathrm{o}}\right]\right\}$
+$+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (51)
+$\hat{n}_{z\left[0\right]\mathrm{qd}}\approx$ $\left[f\left(\theta_{\mathrm{o}}-\alpha;\chi\right)\right]^{-1/2}\left\{2\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\theta_{\mathrm{o}}-\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\theta_{\mathrm{o}}~{}_{~{}_{~{}_{~{}_{~}}}}^{~{}^{~{}^{~{}^{~}}}}\right.$
+$\left.+\hskip 1.0pt2\chi\frac{r_{\ast}}{r_{\mathrm{o}}}\left[\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)-1\right)\cos\theta_{\mathrm{o}}-2\cos\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\theta_{\mathrm{o}}\right]\right\}$
+$+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (52)
+and
+$\hat{n}_{y\left[0\right]\mathrm{qd}}\approx$ $\frac{\Omega_{\mathrm{p}}}{c}\left\{\left[f\left(\theta_{\mathrm{o}}-\alpha;\chi\right)\right]^{-1/2}\sin\alpha\left(1+4\chi\frac{r_{\ast}}{r_{\mathrm{o}}}\cos\left(\theta_{\mathrm{o}}-\alpha\right)\right)s+r_{\mathrm{o}}\sin\theta_{\mathrm{o}}\right\}$
+$+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (53)
+with $f\left(\theta_{\mathrm{o}}-\alpha;\chi\right)$ referring to Eq.
+(49) for $\theta=\theta_{\mathrm{o}}$ and $\phi=0$.
+In association with $\mathbf{\hat{n}}_{\left[0\right]\mathrm{qd}}$, the orthogonal pair of classical mode polarization vectors are determined as
+$\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]\mathrm{qd}}$ $=\hat{n}_{z\left[0\right]\mathrm{qd}}\mathbf{e}_{x}+\hat{n}_{y\left[0\right]\mathrm{qd}}\mathbf{e}_{y}-\hat{n}_{x\left[0\right]\mathrm{qd}}\mathbf{e}_{z},$ (54)
+$\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]\mathrm{qd}}$ $=-\left(\hat{n}_{x\left[0\right]\mathrm{qd}}+\hat{n}_{z\left[0\right]\mathrm{qd}}\right)\hat{n}_{y\left[0\right]\mathrm{qd}}\mathbf{e}_{x}+\mathbf{e}_{y}+\left(\hat{n}_{x\left[0\right]\mathrm{qd}}-\hat{n}_{z\left[0\right]\mathrm{qd}}\right)\hat{n}_{y\left[0\right]\mathrm{qd}}\mathbf{e}_{z},$ (55)
+such that the three vectors, $\mathbf{\hat{n}}_{\left[0\right]\mathrm{qd}}$, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]\mathrm{qd}}$ and $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]\mathrm{qd}}$ form an orthonormal basis.
+In addition, due to (46) and (50), the angle between the photon trajectory and the local magnetic field line, as defined by Eq. (3) shall be modified. Then we have
+$\cos\vartheta_{\mathrm{qd}}$ $\approx\left[f\left(\theta_{\mathrm{o}}-\alpha;\chi\right)\right]^{-1/2}\left[f\left(\theta-\alpha;\chi\right)\right]^{-1/2}$
+×$\hskip 12.0pt\times\left[4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)+\chi g_{1}\left(r,\theta\right)+\chi^{2}g_{2}\left(r,\theta\right)\right]$
+$\hskip 12.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (56)
+where
+$g_{1}\left(r,\theta\right)$ $=4\frac{r_{\ast}}{r_{\mathrm{o}}}\left[\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)-1\right)\cos\left(\theta-\alpha\right)+\cos\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta-\alpha\right)\right]$
+$\hskip 12.0pt+4\frac{r_{\ast}}{r}\left[\left(3\cos^{2}\left(\theta-\alpha\right)-1\right)\cos\left(\theta_{\mathrm{o}}-\alpha\right)+\cos\left(\theta-\alpha\right)\sin\left(\theta-\alpha\right)\sin\left(\theta_{\mathrm{o}}-\alpha\right)\right],$ (57)
+$g_{2}\left(r,\theta\right)$ $=4\frac{r_{\ast}^{2}}{r_{\mathrm{o}}r}\left[\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)-1\right)\left(3\cos^{2}\left(\theta-\alpha\right)-1\right)\right.$
+$\hskip 6.0pt\left.{}_{\,}^{\,}+4\cos\left(\theta_{\mathrm{o}}-\alpha\right)\sin\left(\theta_{\mathrm{o}}-\alpha\right)\cos\left(\theta-\alpha\right)\sin\left(\theta-\alpha\right)\right].$ (58)
+Now, having $\mathbf{B}_{\mathrm{qd}}$, $\mathbf{\hat{n}}_{\left[0\right]\mathrm{qd}}$, $\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]\mathrm{qd}}$, $\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]\mathrm{qd}}$ and $\cos\vartheta_{\mathrm{qd}}$ at hand, as given by Eqs.
+(46), (50), (54), (55) and (56) above, respectively, we modify the evolution equations of the Stokes vector (1) to
+$\frac{\mathrm{d}\mathbf{S}}{\mathrm{d}s}=k\mathbf{\hat{\Omega}}_{\mathrm{qd}}\times\mathbf{S},$ (59)
+where
+$\mathbf{\hat{\Omega}}_{\mathrm{qd}}\equiv\frac{\alpha_{\mathrm{e}}}{30\pi}\left(B_{\mathrm{qd}}/B_{\mathrm{c}}\right)^{2}\sin^{2}\vartheta_{\mathrm{qd}}\left(\mathcal{E}_{\mathrm{I}\mathrm{qd}}^{2}-\mathcal{E}_{\mathrm{II}\mathrm{qd}}^{2},2\mathcal{E}_{\mathrm{I}\mathrm{qd}}\mathcal{E}_{\mathrm{II\mathrm{qd}}},0\right).$ (60)
+Here $\sin\vartheta_{\mathrm{qd}}$ is defined via (56), and
+$\mathcal{E}_{\mathrm{I}\mathrm{qd}}=-\mathbf{\hat{B}_{\mathrm{qd}}\cdot}\left(\mathbf{\hat{n}}_{\left[0\right]\mathrm{qd}}\times\boldsymbol{\varepsilon}_{\mathrm{I}\left[0\right]\mathrm{qd}}\right)$
+$\approx\left[f\left(\theta_{\mathrm{o}}-\alpha;\chi\right)\right]^{-1/2}\left[f\left(\theta-\alpha;\chi\right)\right]^{-1/2}$
+×$\hskip 8.0pt\times\left\{\left[2\cos\left(\theta_{\mathrm{o}}-\alpha\right)+2\chi\frac{r_{\ast}}{r_{\mathrm{o}}}\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)-1\right)\right]\right.$
+×$\hskip 24.0pt\times\left[2\cos\left(\theta-\alpha\right)+2\chi\frac{r_{\ast}}{r}\left(3\cos^{2}\left(\theta-\alpha\right)-1\right)\right]$
+$\hskip 24.0pt\left.+\sin\left(\theta_{\mathrm{o}}-\alpha\right)\left[1+4\chi\frac{r_{\ast}}{r_{\mathrm{o}}}\cos\left(\theta_{\mathrm{o}}-\alpha\right)\right]\sin\left(\theta-\alpha\right)\left[1+4\chi\frac{r_{\ast}}{r}\cos\left(\theta-\alpha\right)\right]\right\}\hat{n}_{y\left[0\right]\mathrm{qd}}$
+$\hskip 8.0pt-\mathcal{E}_{\mathrm{II}\mathrm{qd}}\hat{n}_{y\left[0\right]\mathrm{qd}}-\frac{\Omega_{\mathrm{p}}\sin\alpha}{c}\left[f\left(\theta-\alpha;\chi\right)\right]^{-1/2}\left[1+4\chi\frac{r_{\ast}}{r}\cos\left(\theta-\alpha\right)\right]s$
+$\hskip 8.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (61)
+$\mathcal{E}_{\mathrm{II}\mathrm{qd}}=-\mathbf{\hat{B}_{\mathrm{qd}}\cdot}\left(\mathbf{\hat{n}}_{\left[0\right]\mathrm{qd}}\times\boldsymbol{\varepsilon}_{\mathrm{II}\left[0\right]\mathrm{qd}}\right)$
+$\approx\left[f\left(\theta_{\mathrm{o}}-\alpha;\chi\right)\right]^{-1/2}\left[f\left(\theta-\alpha;\chi\right)\right]^{-1/2}$
+×$\hskip 8.0pt\times\left\{\left[2\cos\left(\theta-\alpha\right)+2\chi\frac{r_{\ast}}{r}\left(3\cos^{2}\left(\theta-\alpha\right)-1\right)\right]\sin\left(\theta_{\mathrm{o}}-\alpha\right)\left[1+4\chi\frac{r_{\ast}}{r_{\mathrm{o}}}\cos\left(\theta_{\mathrm{o}}-\alpha\right)\right]\right.$
+$\hskip 24.0pt\left.-\left[2\cos\left(\theta_{\mathrm{o}}-\alpha\right)+2\chi\frac{r_{\ast}}{r_{\mathrm{o}}}\left(3\cos^{2}\left(\theta_{\mathrm{o}}-\alpha\right)-1\right)\right]\sin\left(\theta-\alpha\right)\left[1+4\chi\frac{r_{\ast}}{r}\cos\left(\theta-\alpha\right)\right]\right\}$
+$\hskip 8.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (62)
+with $f\left(\theta-\alpha;\chi\right)$ referring to Eq. (49) for $\phi=0$.
+3.2 Solving the evolution equations
+From Eq. (60) one can write out the non-zero components of the birefringent vector:
+$\hat{\Omega}_{1\mathrm{qd}}$ $\approx-\eta B^{2}_{\mathrm{qd}}\sin^{2}\vartheta_{\mathrm{qd}}\mathcal{E}^{2}_{\mathrm{II}\mathrm{qd}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (63)
+$\hat{\Omega}_{2\mathrm{qd}}$ $\approx 2\eta B^{2}_{\mathrm{qd}}\sin^{2}\vartheta_{\mathrm{qd}}\mathcal{E}_{\mathrm{I}\mathrm{qd}}\mathcal{E}_{\mathrm{II}\mathrm{qd}}+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (64)
+where $\eta\equiv\alpha_{\mathrm{e}}/\left(30\pi B_{\mathrm{c}}^{2}\right)$, and $\mathcal{E}_{\mathrm{I}\mathrm{qd}}$ and $\mathcal{E}_{\mathrm{II}\mathrm{qd}}$ refer to Eqs.
+(61) and (62), respectively, and $\sin\vartheta_{\mathrm{qd}}$ is defined via Eq.
+(56), while
+$B_{\mathrm{qd}}=\frac{B_{\mathrm{max}}r_{*}^{3}\,\left[f\left(\theta-\alpha;\chi\right)\right]^{1/2}}{2\left(1+2\chi\right)\left(x^{2}+z^{2}\right)^{3/2}}$ (65)
+due to Eq.
+(46), with $B_{\mathrm{max}}$ being the maximum magnetic field intensity at the polar cap, $r_{*}$ being the neutron star radius and $f\left(\theta-\alpha;\chi\right)$ referring to Eq.
+(49) for $\phi=0$.
+Having Eqs.
+(63) and (64) at hand, we solve a system of differential equations, i.e., the evolution equations of the Stokes vector, written out in component form from (59):
+$\dot{S}_{1}\left(s\right)$ $=k\hat{\Omega}_{2\mathrm{qd}}\left(s\right)S_{3}\left(s\right),$ (66)
+$\dot{S}_{2}\left(s\right)$ $=-k\hat{\Omega}_{1\mathrm{qd}}\left(s\right)S_{3}\left(s\right),$ (67)
+$\dot{S}_{3}\left(s\right)$ $=k\left[\hat{\Omega}_{1\mathrm{qd}}\left(s\right)S_{2}\left(s\right)-\hat{\Omega}_{2\mathrm{qd}}\left(s\right)S_{1}\left(s\right)\right].$ (68)
+3.2.1 Examples
+We consider again X-ray emissions from the same three RPPs as in Sect.
+2.2.1: ($\mathrm{i}^{\prime}$) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), ($\mathrm{ii}^{\prime}$) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), ($\mathrm{iii}^{\prime}$) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$).
+Again, for all three, we assume $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$).
+However, unlike the dipole pulsars as in Sect.
+2.2.1, quadrudipole pulsars can be modeled by setting the value of an arbitrary parameter $\chi$ to determine the ratio between the magnetic dipole and quadrupole moments; we choose two values, $\chi=1.585$ and $0.85$ to model each RPP in our analysis here.
+[^14] The values $\chi=1.585$ and $0.85$ have been determined such that they optimize $\hat{\Omega}_{1\mathrm{qd}}\left(s;\chi\right)\leq 0$ and $\hat{\Omega}_{2\mathrm{qd}}\left(s;\chi\right)\geq 0$, respectively.
+That is to say, the optimal values of $\hat{\Omega}_{1\mathrm{qd}\left(\mathrm{min}\right)}$ and $\hat{\Omega}_{2\mathrm{qd}\left(\mathrm{max}\right)}$ are found at $\left(s\approx 8.2444\times 10^{5}\,\mathrm{cm};\chi\approx 1.585\right)$ and $\left(s\approx 6.3588\times 10^{5}\,\mathrm{cm};\chi\approx 0.85\right)$, respectively, which are determined from $\partial\hat{\Omega}_{1\mathrm{qd}}\left(s;\chi\right)/\partial s=\partial\hat{\Omega}_{1\mathrm{qd}}\left(s;\chi\right)/\partial\chi=0$ and $\partial\hat{\Omega}_{2\mathrm{qd}}\left(s;\chi\right)/\partial s=\partial\hat{\Omega}_{2\mathrm{qd}}\left(s;\chi\right)/\partial\chi=0$, respectively.
+In consideration of Eqs.
+(66)-(68), these values will maximize the effects of birefringence on our evolution system.
+In Figs.
+10, 11 and 12 are plotted our numerical solutions of Eqs.
+(66)-(68) for the Stokes vectors in Examples ($\mathrm{i}^{\prime}$), ($\mathrm{ii}^{\prime}$) and ($\mathrm{iii}^{\prime}$), respectively.
+Also, in Fig.
+13 we present the solutions as represented on the Poincaré sphere for $\chi=1.585$ only; there is little difference in the representations of solutions between $\chi=1.585$ and $0.85$ cases.
+The solutions represented by the three magenta loci and the three light blue loci in Fig.
+13 correspond to Figs.
+10a, 11a, 12a and Figs.
+10c, 11c, 12c, respectively.
+In Fig.
+10 for Example ($\mathrm{i}^{\prime}$), where $B_{\mathrm{max}}$ is relatively weak among the three RPPs, general patterns of polarization evolution are shown to be nearly the same for $\chi=1.585$ and $0.85$, and to be almost the same even to the dipole case as given by Fig.
+3 for Example (i) in Sect.
+2.2.1; it is also confirmed by comparing the representations in Figs.
+6a and 13a.
+In contrast, in Fig.
+11 for Example ($\mathrm{ii}^{\prime}$), where $B_{\mathrm{max}}$ is intermediate among the three RPPs, general patterns of polarization evolution are shown to be nearly the same for $\chi=1.585$ and $0.85$, but to be noticeably different from the dipole case as given by Fig.
+4 for Example (ii) in Sect.
+2.2.1; comparing the representations in Figs.
+6b and 13b, the number of cycles appears to become nearly doubled (from half a cycle to one cycle).
+However, in Fig.
+12 for Example ($\mathrm{iii}^{\prime}$), where $B_{\mathrm{max}}$ is the strongest among the three RPPs, general patterns of polarization evolution are shown to be similar for $\chi=1.585$ and $0.85$, with noticeable differences in phase, but to be significantly different from the dipole case as given by Fig.
+5 for Example (iii) in Sect.
+2.2.1; the plots appear to be much denser in the early part of evolution in Fig.
+12 than in Fig.
+5, as the polarization states oscillate much more frequently in the quadrudipole field than in the dipole field, which can also be confirmed by comparing the representations in Figs.
+6c and 13c.
+(a) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$ and $\chi=1.585$
+(b) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$ and $\chi=0.85$
+(c) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$ and $\chi=1.585$
+(d) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$ and $\chi=0.85$
+Figure 10: For Example ($\mathrm{i}^{\prime}$): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the quadrudipole pulsar with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$).
+(a) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$ and $\chi=1.585$
+(b) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$ and $\chi=0.85$
+(c) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$ and $\chi=1.585$
+(d) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$ and $\chi=0.85$
+Figure 11: For Example ($\mathrm{ii}^{\prime}$): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the quadrudipole pulsar with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$).
+(a) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$ and $\chi=1.585$
+(b) With $\mathbf{S}\left(0\right)=\left(1,0,0\right)$ and $\chi=0.85$
+(c) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$ and $\chi=1.585$
+(d) With $\mathbf{S}\left(0\right)=\left(0.8,0,0.6\right)$ and $\chi=0.85$
+Figure 12: For Example ($\mathrm{iii}^{\prime}$): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the quadrudipole pulsar with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$).
+(a) Example ($\mathrm{i}^{\prime}$)
+(b) Example ($\mathrm{ii}^{\prime}$)
+(c) Example ($\mathrm{iii}^{\prime}$)
+Figure 13: Representations of the Stokes vectors from Examples ($\mathrm{i}^{\prime}$)-($\mathrm{iii}^{\prime}$) for $\chi=1.585$ on the Poincaré sphere.
+The loci imply patterns of the polarization evolution in terms of oscillation.
 4 Conclusions and discussion
-We have investigated the evolution of polarization states under the quantum refraction effects combined with the frequency dependence of pulsar emission.
-To this end, we have solved a system of evolution equations of the Stokes vector given by (1) (or by (23)-(25)) for three examples of RPPs at a fixed frequency for specific emissions (e.g., X-rays as in Sections 2.2.1 and 2.2.2).
-Our main results are presented in Figs.
-3-6, from numerical solutions and in part from perturbative solutions.
-Also, we have approximated the birefringent vector by some models as in Fig.
-7 to solve the evolution equations analytically, and obtained the results as presented in Fig.
-8, from approximate analytical solutions.
-It is noteworthy that at a fixed frequency of emission the evolution of the Stokes vector largely exhibits three different patterns, depending on the magnitudes of the birefringent vector, in which the magnetic field strength is a dominant factor: (i) fractionally oscillatory - adiabatic/monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviours.
+We have investigated the evolution of polarization states of pulsar emission under the quantum refraction effects, combined with the dependence on the emission frequency, for both dipole and quadrudipole pulsar models.
+To this end, we have solved a system of evolution equations of the Stokes vector given by (1) (or by (23)-(25)) and (59) (or by (66)-(68)) in the dipole and quadrudipole cases, respectively, for three examples of RPPs at a fixed frequency for specific emissions (e.g., X-rays as in sections 2.2.1, 2.2.2 and 3.2.1).
+Our main results are presented by Figs.
+3-5 and 10-12 in the dipole and quadrudipole cases, respectively, from numerical solutions and in part from perturbative solutions.
+Also, we have replaced the birefringent vector with some approximate models as in Fig.
+7 to solve the evolution equations analytically in the dipole case, and obtained the results as presented by Fig.
+8.
+It is noteworthy that at a fixed frequency of emission the evolution of the Stokes vector largely exhibits three different patterns, depending on the magnitudes of the birefringent vector, in which the magnetic field strength is a dominant factor: (i) fractionally oscillatory - monotonic, or (ii) half-oscillatory, or (iii) highly oscillatory behaviors.
 These features are shown by the numerical solutions in Figs.
@@
 8.
-This study is centred on solving the evolution equations for polarization states (1), wherein the birefringent vector that contains all information about the quantum refraction effects, coupled to the frequency of pulsar emission, acts on the Stokes vector; the evolution results from the combination of the quantum refraction effects and the frequency dependence of the emission.
-This is a major difference from our previous work (Kim et al.
-(2024)), wherein the same effects are independent of the emission frequency; the work solely focuses on the quantum refraction effects on the propagation and polarization of a photon from pulsar emission, with no reference to other properties of the emission, such as the frequency.
+In addition, we have examined how the aforementioned features change in the quadrudipole case.
+This study is centered on solving the evolution equations for polarization states (1), wherein the birefringent vector that contains all information about the quantum refraction effects, coupled to the frequency of pulsar emission, acts on the Stokes vector; the evolution results from the combination of the quantum refraction effects and the frequency dependence of the emission.
+This is a major difference from our previous work 10.1093/mnras/stae1304 , wherein the same effects have no connection to the emission frequency; the work solely focuses on the quantum refraction effects on the propagation and polarization vectors in pulsar emission, with no reference to other properties, such as the emission frequency.
 In this regard, it is worthwhile to draw comparison between the two quantities, the polarization vector and the Stokes vector, both of which are used to describe polarization states.
-The polarization vector is defined directly from the radiative electric field vector (i.e.
-the unit electric field vector), and it is parallel-transported along the the propagation vector; usually, we consider such two vectors orthogonal to each other and to the propagation vector to define an orthonormal basis consisting of the three vectors.
-In contrast, the Stokes vector is defined from Stokes parameters which are built out of the radiative electric field vector (Rybicki & Lightman (1991)).
+The polarization vector is defined directly from the radiative electric field vector (i.e., the unit electric field vector), and it is parallel-transported along the the propagation vector; usually, we consider such two vectors orthogonal to each other and to the propagation vector to define an orthonormal basis consisting of the three vectors.
+In contrast, the Stokes vector is defined from Stokes parameters which are built out of the radiative electric field vector Rybicki:847173 .
 The representation of the Stokes vector is abstract in the sense that it is a vector defined on the Poincaré sphere.
-The Stokes vector is not parallel-transported along the the propagation vector, but can still be defined along the propagation vector as the two polarization vectors move along it; hence, it can be parametrised by $s$ to represent polarization states along the photon trajectory.
-However, the Stokes vector has a crucial advantage over the polarization vector in representing polarization states in some astrophysical studies like this: it can be directly estimated from polarimetric measurements and accommodate depolarization effects due to incomplete coherence and random processes during the photon propagation (Scully & Zubairy (1997)).
-With regard to the adiabatic evolution condition as mentioned in Heyl & Shaviv (2000), Heyl et al.
-(2003), and Heyl & Caiazzo (2018), we have carefully examined our results presented in Figs.
-3-5 to see what interpretations the condition leads to.
-Solving the condition $\left|k\mathbf{\hat{\Omega}}\left(\mathrm{d}\ln|k\mathbf{\hat{\Omega}}|/\mathrm{d}s\right)^{-1}\right|\gtrsim 0.05$ for $s$ (note that our $k\mathbf{\hat{\Omega}}$ is equivalent to the birefringent vector as defined in the references above and that we have set the condition value to $0.05$ rather than $0.5$ as in the references) yields $s_{\mathrm{PL}1}\lesssim s\lesssim s_{\mathrm{PL}2}$, where $s_{\mathrm{PL}1[2]}$ refers to the lower [upper] bound for the ‘polarization limiting’ distance as measured from the emission point.
-Using this, we have checked out the following: (1) for $6.2\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 1.8\times 10^{6}\,\mathrm{cm}$ in Fig.
-3, (2) for $3.1\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 6.2\times 10^{6}\,\mathrm{cm}$ in Fig.
-4, (3) for $1.1\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 1.9\times 10^{7}\,\mathrm{cm}$ in Fig.
-5, our Stokes vector evolves evidently; otherwise, it freezes.
+The Stokes vector is not parallel-transported along the the propagation vector, but can still be defined along the propagation vector as the two polarization vectors move along it; hence, it can be parameterized by $s$ to represent polarization states along the photon trajectory.
+However, the Stokes vector has a crucial advantage over the polarization vector in representing polarization states in some astrophysical studies like this: it can be directly estimated from polarimetric measurements and accommodate depolarization effects due to incomplete coherence and random processes during the photon propagation Scully1997 .
+Our results in this study may be of some significance for the currently operating and planned X-ray space telescopes: Imaging X-ray Polarimetry Explorer (IXPE) IXPE , X-ray Polarimeter Satellite (XPoSat) XPoSat , the enhanced X-ray Timing and Polarimetry mission (eXTP) Santangelo_2019 and the Compton Telescope project Wadiasingh2019Magnetars .
+These telescopes measure the polarization of the X-rays from energetic compact objects such as magnetars and black holes to unveil their geometry and physical environment in detail.
+Several magnetars observed by IXPE have been estimated to have overcritical field strengths Taverna2022Polarized ; Turolla2024IXPE .
+Furthermore, a recent measurement of the X-ray polarization of the magnetar 4U 0142+61 has shown that the polarization degree and angle change as a function of X-ray energy, the interpretation of which has led to two competing scenarios about the X-ray emission of the magnetar Taverna2022Polarized ; Lai2023IXPE .
+In fact, a full analysis of the polarimetry data would require a physical model that comprehensively incorporates the properties regarding the polarization of surface emission, the photon propagation through magnetized plasmas, birefringence due to a magnetized quantum vacuum, and gravitational effects on photon propagation Taverna2015Polarization ; Caiazzo2022Probing .
+In this study, we have focused on vacuum birefringence as it is one of the most significant phenomenological issues to be tested by the X-ray polarimetry in practice.
+When the magnetic field is sufficiently strong and slowly varying, the polarization states evolve due to vacuum birefringence; that is, the Stokes vector components change during the photon propagation within the so-called polarization-limiting radius Heyl2000 ; Heyl2003highenergy , which can be several to a couple dozen neutron-star radii, depending on the magnetic field strength at the surface of a neutron star and the emission frequency.
+However, beyond the polarization-limiting radius, the polarization states ‘freeze’, remaining the same until finally being observed through polarimetry Heyl2003highenergy ; Caiazzo2022Probing .
+The new features of the polarization evolution presented in our study, such as the three different oscillatory patterns of the Stokes vectors and the effects of the possible contributions from the multipolar components may all closely concern the observation through the X-ray polarimetry, and therefore should be taken into proper consideration for a more accurate model for pulsar emission.
 Effects of gravitation have not been considered in this study.
@@
 In this regard, inclusion of the gravitational effects will involve non-trivial and immense analyses, and therefore shall be conducted for a long-term plan in our future studies.
-Acknowledgements
-D.-H.K was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1I1A1A01054781).
+[…]
+Acknowledgments D.-H.K was supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1I1A1A01054781).
 C.M.K.
@@
 was also in part supported by National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2019R1I1A3A01063183).
-Data Availability
-The inclusion of a Data Availability Statement is a requirement for articles published in MNRAS.
-Data Availability Statements provide a standardised format for readers to understand the availability of data underlying the research results described in the article.
-The statement may refer to original data generated in the course of the study or to third-party data analysed in the article.
-The statement should describe and provide means of access, where possible, by linking to the data or providing the required accession numbers for the relevant databases or DOIs.
 Appendix A The classical Stokes vector
@@
 where $t_{\mathrm{ret}}\equiv t-r/c$ is the retarded time, $\boldsymbol{\xi}$ represents the particle’s trajectory, $\mathbf{n}$ is the propagation direction of the radiation, and an over-dot ${\dot{}}$ denotes differentiation with respect to $t$.
-In a suitably chosen Cartesian frame, by setting $\boldsymbol{\xi}\left(t_{\mathrm{ret}}\right)=\rho\left(\sin\left(\beta ct_{\mathrm{ret}}/\rho\right),0,\cos\left(\beta ct_{\mathrm{ret}}/\rho\right)\right)$, with $\rho$ being the radius of curvature of the particle’s trajectory, and $\mathbf{n}=\left(\cos\varphi,\sin\varphi,0\right)$, with $\varphi$ being the angle measured from the $x$-axis to the plane of the particle’s motion, we can construct a simple toy model for pulse profiles of pulsar curvature emission as described below (Kim & Trippe (2021)).
-One can express Stokes parameters out of the radiation field (46), which describe polarization properties of the curvature radiation (Gil & Snakowski (1990)):
+In a suitably chosen Cartesian frame, by setting $\boldsymbol{\xi}\left(t_{\mathrm{ret}}\right)=\rho\left(\sin\left(\beta ct_{\mathrm{ret}}/\rho\right),0,\cos\left(\beta ct_{\mathrm{ret}}/\rho\right)\right)$, with $\rho$ being the radius of curvature of the particle’s trajectory, and $\mathbf{n}=\left(\cos\varphi,\sin\varphi,0\right)$, with $\varphi$ being the angle measured from the $x$-axis to the plane of the particle’s motion, we can construct a simple toy model for pulse profiles of pulsar curvature emission as described below Kim2021Generala .
+One can express Stokes parameters out of the radiation field (69), which describe its polarization properties Gil1990Curvature :
 $I=\tilde{E}_{\parallel}^{\ast}\tilde{E}_{\parallel}+\tilde{E}_{\perp}^{\ast}\tilde{E}_{\perp}$
@@
 where $\tilde{E}_{\parallel}$ and $\tilde{E}_{\perp}$ denote the components of the Fourier transform $\mathbf{\tilde{E}}\left(\omega\right)=\tilde{E}_{\parallel}\left(\omega\right)\mathbf{e}_{z}+\tilde{E}_{\perp}\left(\omega\right)\,\mathbf{e}_{y}=\int_{-\infty}^{\infty}\mathbf{E}\left(t\right)\exp\left(\mathrm{i}\omega t\right)\,\mathrm{d}t$, expressed in the Cartesian frame, and means the complex conjugate, and $\mathcal{E}_{\mathrm{o}}=q\beta/\left(2\sqrt{3}\pi^{2}r\rho\right)$, and $\delta\equiv\gamma^{-1}=\left(1-\beta^{2}\right)^{1/2}\ll 1$ is the half-angle of the beam emission, and $\mathrm{K}_{1/3}$ and $\mathrm{K}_{2/3}$ denote the modified Bessel functions of the second kind.
-With regard to the polarization state of the curvature radiation, $I$ is a measure of the total intensity, $Q$ and $U$ jointly describe the linear polarization, and $V$ describes the circular polarization.
+With regard to the polarization state of the radiation field, $I$ is a measure of the total intensity, $Q$ and $U$ jointly describe the linear polarization, and $V$ describes the circular polarization.
 These parameters can be plotted as functions of the phase angle $\varphi$, where $\varphi\leq\delta\ll 5^{\circ}$ usually, to simulate the pulse profiles of pulsar emission theoretically.
@@
 14 is plotted the classical Stokes vector against the phase angle $\varphi$, where we have set, for example, $\delta\approx 10^{-8}$, $\rho\approx 10^{8}\,\mathrm{cm}$, $\beta\approx 1$ and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ to model pulse profiles of X-ray pulsar emission.
-Here the initial values for the Stokes vector, $\mathbf{S}\left(\varphi=0\right)=\left(1,0,0\right)$ and $\mathbf{S}\left(\varphi\approx-1.16\times 10^{-6}\,\mathrm{rad}\right)=\left(0.8,0,0.6\right)$, as in the examples given in Section 2.2.1, are marked by solid circles and solid boxes, respectively.
-Figure 9: The classical Stokes vector $\mathbf{S}\left(\varphi\right)=\left(S_{1}\left(\varphi\right),S_{2}\left(\varphi\right),S_{3}\left(\varphi\right)\right)$ plotted against the phase angle $\varphi$.
+Here the initial values for the Stokes vector, $\mathbf{S}\left(\varphi=0\right)=\left(1,0,0\right)$ and $\mathbf{S}\left(\varphi\approx-1.16\times 10^{-6}\,\mathrm{rad}\right)=\left(0.8,0,0.6\right)$, as in the examples given in Sect.
+2.2.1, are marked by solid circles and solid boxes, respectively.
+Figure 14: The classical Stokes vector $\mathbf{S}_{\left(\mathrm{c}\right)}\left(\varphi\right)=\left(S_{1\left(\mathrm{c}\right)}\left(\varphi\right),S_{2\left(\mathrm{c}\right)}\left(\varphi\right),S_{3\left(\mathrm{c}\right)}\left(\varphi\right)\right)$ plotted against the phase angle $\varphi$.
 Its initial values $\mathbf{S}\left(\varphi=0\right)=\left(1,0,0\right)$ and $\mathbf{S}\left(\varphi\approx-1.16\times 10^{-6}\,\mathrm{rad}\right)=\left(0.8,0,0.6\right)$ are marked by solid circles and solid boxes, respectively.
 Appendix B Approximate analytical solutions to evolution equations
-Substituting equation (30) into equations (23)-(25), the evolution equations can be reduced as follows:
+Substituting Eq. (30) into Eqs. (23)-(25), the evolution equations can be reduced as follows:
 For $0\leq s\leq 20r_{*}$,
@@
 $\ddot{S}_{3}\left(s\right)-\left(\frac{p+1}{ps}-b\right)\dot{S}_{3}\left(s\right)+k^{2}\left(a_{1}^{2}+a_{2}^{2}\right)\left(s^{\frac{p+1}{p}}e^{-bs}\right)^{2}S_{3}\left(s\right)\approx 0.$ (79)
-First, we solve equation (56) for $S_{3}\left(s\right)$, and then using this solution, obtain $S_{1}\left(s\right)$ and $S_{2}\left(s\right)$, by integrating equations (54) and (55), respectively:
+First, we solve Eq. (79) for $S_{3}\left(s\right)$, and then using this solution, obtain $S_{1}\left(s\right)$ and $S_{2}\left(s\right)$, by integrating Eqs. (77) and (78), respectively:
 $S_{3}\left(s\right)$ $\approx S_{\mathrm{o}}\sin\left(\Psi\left(s;p\right)+\delta\right),$ (80)
@@
 $S_{1\mathrm{o}}=\frac{Ca_{1}}{\sqrt{a_{1}^{2}+a_{2}^{2}}},~{}~{}S_{2\mathrm{o}}=\frac{Ca_{2}}{\sqrt{a_{1}^{2}+a_{2}^{2}}},~{}~{}S_{\mathrm{o}}^{2}+C^{2}=1.$ (84)
-Then $S_{\mathrm{o}}$, $C$ and $\delta$ are determined by matching the initial value of the Stokes vector $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)$ with equations (57)-(59) evaluated at $s=0$.
+Then $S_{\mathrm{o}}$, $C$ and $\delta$ are determined by matching the initial value of the Stokes vector $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)$ with Eqs.
+(80)-(82) evaluated at $s=0$.